### v1

# Bitune: Bidirectional Instruction-Tuning

Abstract
We introduce Bitune, a method that improves instruction-tuning of pretrained decoder-only large language models, leading to consistent gains on downstream tasks. Bitune applies both causal and bidirectional attention to the prompt, to obtain a better representation of the query or instruction. We realize this by introducing two sets of parameters, for which we apply parameter-efficient finetuning techniques. These causal and bidirectional features are then combined into a weighted average with trainable coefficients, which is subsequently used to generate new tokens. We demonstrate significant improvements in zero-shot performance on commonsense reasoning, arithmetic, and language understanding tasks, while extensive ablation studies validate the role of each component and demonstrate the method’s agnosticism to different PEFT techniques.
1 Introduction
Large Language models (LLMs) are being used in many practical scenarios where humans interact with them through questions and answers. In applications such as general purpose assistants (OpenAI, 2024a), medical diagnosticians (Thirunavukarasu et al., 2023), game-conversation generation (Cox & Ooi, 2023) or coding-assistants (Roziere et al., 2023), the ability for an LLM to precisely follow an instruction and answer a question are of primary concern.
Correspondingly, Instruction-Tuning (IT)  (Chung et al., 2024; Ouyang et al., 2022a) is the prevailing paradigm for finetuning LLMs after their self-supervised pretraining phase to improve them for such tasks. Here, the model is trained on a dataset comprised of pairs of instructions and corresponding responses. Given the instruction-with-response structure of IT data, the generation of an LLM response can be divided into two phases: first, converting the instruction into Key and Value (KV) embeddings, stored in a KV-cache, which we refer to as instruction features; second, using these features to autoregressively generate an answer. Due to this task’s inherently conditional nature, the instruction features’ effectiveness is crucial for obtaining high-quality model outputs.
In the past, bidirectional attention (Schuster & Paliwal, 1997) has been a key technique for obtaining stronger features for words or tokens. This is because the meaning of a word depends greatly on its context. In particular, for some words in a sentence, the information that comes later might be far more informative for generating a meaningful representation and resolving ambiguities. But with only uni-directional causal attention, where the representation of each word is restricted to depend solely on the words that came before, this cannot be achieved. This is the reason why many previous transformers such as encoder-only BERT (Devlin et al., 2019) and encoder-decoder T5 (Raffel et al., 2020) employed bidirectional attention to improve the encoding of the input and why tasks like text retrieval (Lewis et al., 2020; Li & Li, 2023) and even the latest text-to-image generative models (OpenAI, 2024b; Esser et al., 2024) still rely on this.
However, in the context of LLMs, architectures utilizing bidirectional attention have fallen out of favor, as decoder-only models such as GPT (OpenAI, 2024a) and Llama (AI@Meta, 2024) have focused on and vastly improved the generative performance of language models. These architectures are trained by large volumes of data with next-token prediction, eschewing any look-ahead mechanism for the sake of better autoregressive modeling. As there is simply more unlabeled data available for pretraining, training a decoder-only architecture on unlabeled data, and then finetuning it for tasks with instruction-tuning, is the best modus operandi of today (Wang et al., 2022). However, with this switch to decoder-only architectures, we lost bidirectional attention in the process. As we know this can improve feature representations for instructions, we set out to re-introduce bidirectional attention, such that it can be integrated into pretrained decoder-only LLMs.
Our new method Bitune adds bidirectional attention to decoder-only architectures and combines it with causal attention to generate two sets of instruction features, using two different sets of weights. These features are then integrated, utilizing learnable mixing coefficients, and later used as the KV-cache for response generation. Notably, the autoregressive response generation process remains unaffected by the bidirectional attention and continues to be causal. By realizing these adaptations with parameter-efficient finetuning methods, we introduce only a minimal set of new parameters.
Overall, our contributions are as follows:
  • We propose a novel method, Bitune, that improves the performance of pretrained decoder-only LLMs in question-answering scenarios.
  • We evaluate the method on multiple downstream tasks, showing consistent improvements over the baselines.
  • We conduct an extensive ablation study investigating the necessity of each component of the method, and showing the method’s PEFT-agnosticism.
Figure 1: Overview of Bitune. (a) During the prefilling phase, features are obtained from the prompt using both causal and bidirectional attention in two passes with separate weights. The two sets of keys and values are then combined using a weighted average before being passed to the decoding phase. (b) During the decoding phase, new tokens are generated in the standard way with causal attention, utilizing the features extracted from the instruction in the previous step, along with the features of other generated tokens.
2 Bidirectional Instruction-Tuning
Algorithm 1 Python-like pseudocode of Bitune inference.
# prompt - tensor with tokenized instruction
# theta - tensor with mixing coeff. for each layer
# theta_init - initial value of mixing coefficients
k_c, v_c = model_causal(prompt) # Pass on the instruction for causal features
k_b, v_b = model_bidir(prompt) # Pass to obtain bidirectional features
# Combine both sets of features
alpha = theta.abs() / (theta.abs() + theta_init)
k = k_c * (1 - alpha) + k_b * alpha
v = v_c * (1 - alpha) + v_b * alpha
kv = (k, v)
c_token = SEP # Initialize generation with a predefined token
answer = [c_token]
while c_token != EOS: # Stop generation on the end-of-sequence token
# Get features of current token and logits of next token
k, v, logits = model_causal(c_token, kv)
kv = concat(kv, (k, v)) # Concatenate obtained features with current KV cache
c_token = get_token(logits) # Determine next predicted token based on logits
answer.append(c_token) # Append generated token to the answer
In the instruction-tuning setting (Ouyang et al., 2022b; Zhang et al., 2024), a dataset $\mathcal{D}$ consists of instruction-answer pairs that are used to adapt the model in a supervised fashion. Formally, a dataset of size $N$ can be described as $\mathcal{D}=\{q,a\}^{N}_{i=1}$, where $q$ and $a$ are instructions and answers. The training objective is to model $p(a|q)$ in an autoregressive manner: This means the answer is generated one token at a time, such that token $a_{i}$ at position $i$ has access to all earlier tokens:
$p(a|q)=\Pi^{|a|}_{i=1}p(a_{i}|a_{1},\dots a_{i-1},q),$
where $|a_{i}|$ denotes the length of the answer. Note how compared to the regular language modeling objective, the response is already conditional (on the instruction $q$) even for the first generated token.
This naturally leads response-generation to be divided into two phases: prefilling and decoding. During the prefilling phase, the entire instruction – also often called a prompt – is processed concurrently to generate a series of features to be stored. For a Transformer architecture (Vaswani et al., 2017), these features are those of the key and value vectors, which can be stored in a KV-cache to avoid costly recomputations. During the subsequent decoding phase, the model generates output tokens sequentially, one token at a time, based on the KV-cache of the instruction and the already generated tokens.
In this work, we introduce, Bitune, a method to leverage this two-phase process to improve instruction-tuning of language models. In our approach, the model processes the instruction with both causal and bidirectional attention using separate sets of parameters, leading to an enhanced KV-cache that is then used to condition the answer. Figure 1 provides an overview of the method, while Algorithm 1 presents pseudocode for the inference process.
Two Sets of Features.
In Bitune, the model performs two passes on the instruction to obtain two kinds of features for every transformer block. Namely, a set of causal features that the model was originally trained to process and utilize,
$K_{\color[rgb]{1,0,0}c}=X_{\color[rgb]{1,0,0}c}W_{k{\color[rgb]{1,0,0}c}},\quad V_{\color[rgb]{1,0,0}c}=X_{\color[rgb]{1,0,0}c}W_{v{\color[rgb]{1,0,0}c}},$ (1)
and a set of bidirectional features encoding the instruction without the constraints of causal masking,
$K_{\color[rgb]{0,0,1}b}=X_{\color[rgb]{0,0,1}b}W_{k{\color[rgb]{0,0,1}b}},\quad V_{\color[rgb]{0,0,1}b}=X_{\color[rgb]{0,0,1}b}W_{v{\color[rgb]{0,0,1}b}}.$ (2)
To allow the model to learn how to process the causal and bidirectional features differently, we introduce two sets of weights: one for the bidirectional pass on the instruction ($W_{k{\color[rgb]{0,0,1}b}}$, $W_{v{\color[rgb]{0,0,1}b}}$) and another for the causal pass on the instruction, which is also used for the causal generation of answer tokens ($W_{k{\color[rgb]{1,0,0}c}}$, $W_{v{\color[rgb]{1,0,0}c}}$).
In the case of the first block of the model, representations $X_{\color[rgb]{1,0,0}c}$, $X_{\color[rgb]{0,0,1}b}$ are the initial token embeddings. In other cases, they are the output of the preceding block and were processed by different components including the self-attention mechanism, which can be defined as:
$\text{Attention}(Q,K,V,M)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V,$ (3)
where $Q$ are the queries, $M$ is the attention mask, and $d_{k}$ is the dimension of keys and queries.
For the causal pass, the mask $M_{\color[rgb]{1,0,0}c}$ enforces causality by masking future tokens, such that tokens $j$ can only attend to earlier tokens $i\leq j$, while for the bidirectional pass, no masking is applied:
$M_{{{\color[rgb]{1,0,0}c}}}(i,j)=\begin{cases}0&\text{if }i\leq j\\ -\infty&\text{if }i>j\end{cases},\qquad M_{\color[rgb]{0,0,1}b}(i,j)=0.$ (4)
The final KV-cache is obtained by a learnable convex combination of causal and bidirectional features,
$K_{\texttt{Bitune}}=K_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{k})+K_{\color[rgb]{0,0,1}b}\cdot\alpha_{k},\quad V_{\texttt{Bitune}}=V_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{v})+V_{\color[rgb]{0,0,1}b}\cdot\alpha_{v},$ (5)
where $\alpha$ represents the bidirectional-to-causal ratio of features. This ratio is parameterised as
$\alpha_{j}=\frac{|\theta_{j}|}{\theta_{\text{init}}+|\theta_{j}|},\quad j\in\{k,v\}$ (6)
where $\theta_{j}$ is a learnable mixing coefficient per transformer block, and $\theta_{\text{init}}$ is a hyperparameter defining the initial value of $\theta_{j}$. The mixing coefficients are learnable to allow each block to independently adjust the balance between bidirectional and causal features throughout the training.
Parameter Efficient Fine-tuning
Note that the components of the model, other than the key and value projections, can have their own separate sets of weights as well. In the case of full finetuning, this approach would require an additional set of full weights, which is impractical for large models.
Instead, we adapt our model to use bidirectional attention by using parameter-efficient finetuning methods. These introduce only a fraction of trainable parameters, making it viable to have two modified variants of the model within a single forward pass. In the default configuration of our method, we utilize the Low-Rank Adaptation (LoRA) of Hu et al. (2022) to adapt the model. However, our method is PEFT-agnostic and can also utilize different methods for updating the weights, as demonstrated in our ablations section.
In summary, Bitune is a method for improved instruction-tuning that learns a combination of causal and bidirectional features by applying two sets of lightweight PEFT methods to a pretrained model.
3 Experiments
Table 1: Zero-shot results after instruction-tuning on the UltraFeedback dataset. We compare Bitune to the performance of the original model, the model finetuned with LoRA, and with LoRA16 using two times higher rank to match the parameter count of our method. Bitune significantly outperforms the baselines on almost all tasks for all models.
PIQA ARC CSQA SIQA MMLU Avg.
Model Method
Gemma-2B Pretrained 57.5 36.9 35.5 38.2 34.0 40.4
LoRA 66.7 43.4 42.3 44.3 31.7 45.7
[…]
Bitune 69.6 47.5 46.9 49.5 35.3 49.7
Gemma-7B Pretrained 73.1 78.3 62.0 64.7 59.0 67.4
LoRA 84.2 79.2 68.5 71.9 55.3 71.8
[…]
Bitune 83.6 80.1 69.2 72.7 53.8 71.9
Llama2-7B Pretrained 59.2 38.1 32.6 45.1 36.0 42.2
LoRA 69.5 49.9 45.3 57.0 41.1 52.6
[…]
Bitune 70.0 51.1 48.1 59.1 41.9 54.0
Llama3-8B Pretrained 69.0 73.6 65.4 56.8 56.0 64.2
LoRA 81.9 74.5 69.2 69.0 57.6 70.4
[…]
Bitune 84.4 77.4 72.7 70.1 59.0 72.7
Phi-2 Pretrained 70.3 67.3 61.4 65.0 45.4 61.9
LoRA 76.3 66.7 61.6 66.6 48.2 63.9
[…]
Bitune 76.5 67.2 63.0 68.5 48.9 64.8
3.1 Instruction-Tuning
Our core experiments involve training pretrained language models on an instruction-tuning dataset and zero-shot evaluating them on downstream tasks. We evaluate Bitune on multiple models, comparing results to standard finetuning with LoRA, and zero-shot results of pretrained models without finetuning.
Specifically, we use a subset of the cleaned UltraFeedback (Cui et al., 2023) dataset, which contains instructions and corresponding answers generated by various LLMs. From this dataset, we select completions generated by GPT-4 (OpenAI, 2024a), ensuring high-quality responses for training. To fit every model on a single GPU, we filter out samples longer than 512 tokens, which leaves us with around 10,000 samples for training.
We test the method on pretrained decoder-only language models of two different scales of approximately 2 billion and 7 billion parameters. The specific models used in our experiments are: Gemma 2B and 7B (Gemma Team et al., 2024), Llama2 7B (Touvron et al., 2023) and Llama3 8B (AI@Meta, 2024), and Phi-2 (Li et al., 2023), which has 2.7 billion parameters. We use HuggingFace Transformers (Wolf et al., 2020) implementation of these models. We limit our experiments to these specific models due to their popularity and also our limited resources.
[…]
For each model, we tune the learning rate on the LoRA baseline using steps on the approximate logarithmic scale ($1\mathrm{e}{-4}$, $3\mathrm{e}{-4}$, $1\mathrm{e}{-3}$, $3\mathrm{e}{-3}$), and then apply the same rate to the other approaches. Note that this potentially puts our method at a disadvantage compared to the LoRA baseline. All hyperparameters are reported in the Appendix 6.2.
Models are evaluated zero-shot on multiple-choice tasks to assess their performance. For commonsense reasoning, we use the PIQA (Bisk et al., 2020), CommonsenseQA (Talmor et al., 2019), ARC-Challenge (Clark et al., 2018), and SIQA (Sap et al., 2019) datasets, while for language understanding, we use the MMLU (Hendrycks et al., 2021) benchmark. Each task consists of a series of questions, each with multiple choices, where only one answer is correct. As the tasks follow the question-answer pattern, they are compatible with the instruction-tuning setting.
For evaluation, we use the Language Model Evaluation Harness framework (Gao et al., 2023). This framework formats each question using a predefined template, tokenizes the question-choice pairs, runs them through the model, and compares the log-likelihoods of the choices to determine the selected answer. For each model and approach configuration, we conduct experiments using three different random seeds, and average the results.
Models are loaded and trained using bfloat16 precision, except for the mixing part, which operates in the full 32-bit floating-point format. This high level of precision for the mixing of features is important, as minor numerical inaccuracies in the learnable coefficients and intermediate results of the mixing operation may lead to significant deviations in the model’s behavior.
In the decoding phase of the inference with Bitune, to initiate generation, the model requires at least a single token to obtain a set of attention queries, in addition to the keys and values extracted from the instruction. To facilitate this, one can introduce a new learnable <sep> token that would be placed at the beginning of modeled answer, or utilize an existing token. For our experiments, we opted to move the last token of the instruction template to the beginning of the modeled answer. For details on the instruction template used, please refer to the section 6.5 of the Appendix.
Results.
Table 1 shows consistent and significant gains after instruction-tuning with Bitune, with the highest gains seen on the Gemma-2B model, showing a 4 percentage point (pp) improvement over the baseline LoRA and a 9.3 pp improvement over the pretrained model. For the other models, the average gains over baseline finetuning are equal to 1.8, 1.4, and 0.9 pp, for Llama3-8B, Llama2-7B, and Phi-2 respectively.
It is worth noting that the Gemma-7B model shows the lowest average improvement across all tasks, with merely 0.1 pp gain over the baseline finetuning. It is also a single case where the baseline pretrained model achieved the highest score on a task, MMLU, with degraded performance in all fine-tuning approaches. However, this is not an issue with the model’s scale, as significant gains are observed with the Llama2-7B and Llama3-8B models.
3.2 Downstream Task Training
This complementary experiment verifies if Bitune increases the capacity of the model within the narrow scope of a single task. It follows the setup from the instruction-tuning experiments with a few changes. Namely, models are not instruction-tuned but trained separately for each evaluation task using the corresponding training set. We use PIQA, ARC, CSQA, and SIQA introduced earlier, and an additional arithmetic task, GSM8K (Cobbe et al., 2021).
GSM8K differs from the other tasks, where we compare log-likelihoods of predefined answers, as it requires the model to generate a full answer token-by-token, including the intermediate step-by-step reasoning. The final answer follows a specific pattern, making it feasible to extract the answer using methods such as regular expressions as the model learns to adhere to this pattern during training.
Results.
Table 2: Result for the downstream task training. We show accuracy on downstream tasks for the baseline LoRA finetuning and Bitune, averaged over 3 seeds. On average, our method works better than standard LoRA. We see the most significant gains on the GSM8K dataset, but slightly lower results for Gemma-7B and the SIQA task.
PIQA ARC CSQA SIQA GSM8K Avg.
Model Method
Gemma-2B LoRA 81.4 58.0 77.2 77.4 30.2 64.8
Bitune 83.3 60.0 78.3 76.6 33.0 66.2
Gemma-7B LoRA 91.4 84.6 84.4 79.4 59.1 79.8
Bitune 92.1 84.2 84.2 79.4 59.4 79.9
Llama2-7B LoRA 84.4 66.6 81.5 82.7 32.0 69.4
Bitune 84.4 66.9 82.0 81.4 32.9 69.5
Llama3-8B LoRA 90.2 80.7 83.9 83.1 60.4 79.7
Bitune 90.5 81.3 84.1 82.1 63.4 80.3
Phi-2 LoRA 82.8 76.3 78.7 80.3 58.6 75.3
Bitune 83.9 77.0 79.0 80.4 59.2 75.9
Table 2 presents the results, demonstrating improvements when finetuning on the downstream tasks with Bitune, similar to those seen with instruction-tuning. While there are a few cases where the baseline finetuning achieves better results on specific tasks, when considering the average gains, applying our method is beneficial across all models. Most importantly, on the GSM8K dataset, we see consistently high gains, suggesting that our method improves the model’s reasoning ability in generative tasks. The qualitative example for instruction-tuned model in Table 6 also supports this (more examples in Appendix 6.7). Similar to the instruction-tuning results, the highest gains are observed on the Gemma-2B model, while the lowest on the Gemma-7B. This indicates that the effectiveness of our method depends on the specific model used.
3.3 Ablations
We conduct an ablation study on Bitune using the same experimental setup as in the instruction-tuning experiment. For this purpose, two models are used: Gemma-2B and Llama3-8B, representing different size scales and model families.
Component Removal
Table 3: Ablation study on components of Bitune. We report zero-shot accuracy averaged over PIQA, ARC, CSQA, SIQA and MMLU tasks. The components are explained in section 3.3. We see that all ablated variants outperform the LoRA baseline, and combining all components performs the best. Of note, especially bidirectional attention improves results the most.
Avg. Acc.
Model Method Causal Bidir. Mixing Sep. Weights
Gemma-2B LoRA - - - 45.7
Naive Bidir. - - - 47.9
No Mixing - - 48.9
Only Causal - 46.9
Shared Weights - 47.4
Bitune 49.7
Llama3-8B LoRA - - - 70.4
Naive Bidir. - - - 71.9
No Mixing - - 71.5
Only Causal - 71.1
Shared Weights - 72.3
Bitune 72.7
To verify the necessity of each component of the method, we remove selected parts to answer the following questions:
  • Can we simply modify the attention mask to apply bidirectional attention on the prompt, without using separate weights and mixing? - We test this simplest variant, which we refer to as Naive Bidir.
  • Do we need two sets of features? Is it sufficient to obtain bidirectional features from the prompt using different weights than those used for causal answer generation? - We remove the part responsible for generating the set of causal features, and therefore also the mixing component; we refer to this as No Mixing.
  • Are the gains solely from mixing two sets of features generated with different weights, or is bidirectional attention necessary? - Here we keep the attention mask causal to generate both sets of features, which we refer to as Only Causal.
  • Do we need separate weights, or can the same weights be used to generate both causal and bidirectional features? - To answer this question, we do not introduce the second set of weights and use the same LoRA for both passes on the prompt, calling it Shared Weights.
The results, averaged over three seeds and presented in Table 3, indicate that all variants of the method lead to gains over the baseline LoRA finetuning. However, the highest gains are observed in the full variant of Bitune, demonstrating that each component contributes to the method’s effectiveness.
Different PEFT Methods
Table 4: Combining Bitune with different PEFT methods. Performances are averaged over 3 seeds. We can see that our method improves results regardless of the specific PEFT method used.
PIQA ARC CSQA SIQA MMLU Avg.
Model Method
Gemma-2B LoRA 66.7 43.4 42.3 44.3 31.7 45.7
Bitune 69.6 47.5 46.9 49.5 35.3 49.7
DoRA 66.7 43.6 41.9 44.7 31.9 45.8
[…]
IA3 67.2 46.5 45.5 37.6 32.5 45.9
[…]
Llama3-8B LoRA 81.9 74.5 69.2 69.0 57.6 70.4
Bitune 84.4 77.4 72.7 70.1 59.0 72.7
DoRA 82.1 75.4 70.2 69.2 57.7 70.9
[…]
IA3 80.9 75.5 68.3 66.4 58.7 70.0
[…]
To verify the impact of different PEFT methods on the performance of our method, we compare Bitune in combination with the following techniques: LoRA (Hu et al., 2022), that reparametrizes weight updates as a multiplication of two low-rank matrices; DoRA (Liu et al., 2024), which decomposes these weight updates into direction and magnitude; and IA3 (Liu et al., 2022), that instead rescales activations with learnable vectors.
The results are shown in Table 4. We find consistent gains across all three PEFT methods we analyze, with gains ranging from $+1.6\%$ to $+4.0\%$ for averaged accuracies. This demonstrates that Bitune is PEFT-agnostic and can be combined with existing and future innovations in PEFT methods.
Initialization of Mixing Coefficient
The initial value of the mixing coefficient $\theta$ is a hyperparameter in our method. To evaluate its impact on the performance and the training dynamics of the bidirectional-to-causal ratio of features, we conduct experiments on the instruction-tuning setup with the following values: $0.1$, $0.01$, and $0.001$.
Figure 2: Bidirectional-to-causal ratio during training. The ratio is averaged over all layers and shown for different initial values of mixing coefficients for Llama3-8B. The initial value impacts the change of the ratio, with higher values slowing it down, and lower values increasing it.
Figure 3: Ratio across layers. Here we show the final ratio of the model in Fig. 2 across all layers for the K and V values. The utilization of bidirectional attention is spread across all layers.
Table 5(a) demonstrates that the initial value of the mixing coefficient impacts the performance, with $0.01$ being the most optimal value for both models, regardless of their scale. Figure 2 shows that the initial value substantially affects the rate of change of the mixing ratio, with the higher value leading to nearly no change in the ratio, while the lower value results in sharp changes at the very beginning of the training. From Figure 3, we also observe that after training, all layers utilize the bidirectional attention.
Attention Mask of Second Pass
We test another option for the attention mask of the second pass on the instruction. We transpose the causal attention mask, blocking information flow from the past tokens, and allowing from the future tokens - we call it anti-causal attention mask.
Table 5: Ablation of Bitune’s attention. We vary the additional attention mask for processing the instruction besides the causal pass, and evaluate different initial values of mixing coefficient, $\theta_{\text{init}}$.
(a) Different init. values for mixing coefficients.
Model Init. Value Avg. Acc.
Gemma-2B $0.1$ 49.4
$0.01$ 49.7
$0.001$ 47.2
Llama3-8B $0.1$ 72.7
$0.01$ 72.7
$0.001$ 72.3
(b) 2nd pass attention masks for instruction features.
Model Attention Mask Avg. Acc
Gemma-2B Causal 46.9
Anti-causal 46.7
Bidirectional 49.7
Llama3-8B Causal 71.1
Anti-causal 71.4
Bidirectional 72.7
Results shown in Table 5(b) indicate that the instruction has to be processed with full bidirectional attention to achieve the highest gains. Combining causal and anti-causal features independently does not lead to the same high performance.
Table 6: Qualitative results on Llama3-8B. We show a response for LoRA and Bitune.
[GSM8K]     Question: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market?
LoRA Janet makes \$16 - (3 + 4) = \$9 per day at the farmers’ market.
Bitune Janet makes \$2 per fresh duck egg, and she sells the remainder of the eggs she collects each day. If she eats three eggs for breakfast and bakes muffins with four eggs, she has 16 - 3 - 4 = 9 eggs left to sell. So, she makes \$2 x 9 = \$18 every day at the farmers’ market.
4 Related Work
Our approach shares similarities with the concept of "prefix language modeling", which enables a decoder-only model to handle bidirectional context within a prefix (instruction) while maintaining causal generation for the output sequence. The prefix-LM architecture was introduced by Liu et al. (2018) and further explored and popularized by Raffel et al. (2020). In their work on T5, Raffel et al. (2020) pretrained the prefix-LM architecture alongside other architectures, such as encoder-decoder and decoder-only models, demonstrating that prefix-LM outperforms decoder-only models on both training objectives: denoising and language modeling.
The prefix-LM approach has been used in UniLM (Dong et al., 2019), which trains a single transformer on three types of language modeling tasks: unidirectional, bidirectional, and sequence-to-sequence prediction. UniLM employs a shared Transformer network and utilizes specific self-attention masks to control the context that predictions are conditioned on, where the sequence-to-sequence task is equivalent to the prefix-LM approach.
Additionally, UL2 (Tay et al., 2023) introduces a pretraining objective called "Mixture of Denoisers", which combines various denoising strategies, including the prefix-LM approach. Lastly, XLNet Yang et al. (2019) also allows for non-causal word ordering by allowing random permutations to be used with a next-token prediction objective.
All these works focused on the model pretraining. As for the utilization of pretrained causal language models, Springer et al. (2024) show in their work that simply repeating the input to these models improves the quality of token embeddings for text-retrieval. This work addresses the limitation that token embeddings in autoregressive models cannot contain information from tokens appearing later in the input. By repeating the input twice, the early tokens are allowed to encode information about later tokens, thereby improving the quality of the embeddings. Another approach, LLM2Vec (BehnamGhader et al., 2024), demonstrates that pretrained causal LLMs can be effectively converted to BERT-like encoders. It can be done by enabling bidirectional attention, training the model on the objective of masked token prediction, and applying unsupervised contrastive learning.
5 Discussion
Limitations.
During standard instruction-tuning training, the instruction and the answer are processed in a single forward pass. In our method, this processing is explicitly split into phases, extracting instruction features with two passes & answer modeling, increasing both training time and memory usage. As for the inference, these two passes on the instruction, bidirectional and causal, can be processed either in parallel, impacting memory usage, or sequentially, increasing latency.
However, this is a minor limitation in the context of instruction-tuning, since typically smaller datasets are used compared to pretraining, leading to relatively short training times. Furthermore, at inference time the added latency for processing the instruction is negligible, as the bulk of compute is used for the autoregressive answer generation: e.g. only a 0.2s increase from 11.5s to 11.7s for a 200 token response to a 50 token instruction for Llama3-8B (see Appendix 6.4 for more results).
Conclusion.
Bitune is the first method to utilize the instruction-answer structure of IT datasets to propose a new finetuning method that enables bidirectional processing in pretrained decoder-only models. Bitune demonstrates general applicability across different models and scales. It consistently improves performance across commonsense reasoning, arithmetic, and language understanding tasks. We further demonstrate that our method is compatible with different existing PEFT methods and will likely benefit from further developments in this era. As the first work to demonstrate the potential of an instruction-tuning-specific adaptation for decoder LLMs, we believe there is much space for further research into this promising and crucial direction.
Acknowledgements
We acknowledge the use of the National Supercomputer Snellius and the Distributed ASCI Supercomputer 6 (Bal et al., 2016) for essential computational tasks.

### v2

@@
-Bitune: Bidirectional Instruction-Tuning
+Bitune: Leveraging Bidirectional Attention to Improve Decoder-Only LLMs
 Abstract
-We introduce Bitune, a method that improves instruction-tuning of pretrained decoder-only large language models, leading to consistent gains on downstream tasks.
-Bitune applies both causal and bidirectional attention to the prompt, to obtain a better representation of the query or instruction.
-We realize this by introducing two sets of parameters, for which we apply parameter-efficient finetuning techniques.
-These causal and bidirectional features are then combined into a weighted average with trainable coefficients, which is subsequently used to generate new tokens.
-We demonstrate significant improvements in zero-shot performance on commonsense reasoning, arithmetic, and language understanding tasks, while extensive ablation studies validate the role of each component and demonstrate the method’s agnosticism to different PEFT techniques.
+Decoder-only large language models typically rely solely on masked causal attention, which limits their expressiveness by restricting information flow to one direction.
+We propose Bitune, a method that enhances pretrained decoder-only LLMs by incorporating bidirectional attention into prompt processing.
+We evaluate Bitune in instruction-tuning and question-answering settings, showing significant improvements in performance on commonsense reasoning, arithmetic, and language understanding tasks.
+Furthermore, extensive ablation studies validate the role of each component of the method, and demonstrate that Bitune is compatible with various parameter-efficient finetuning techniques and full model finetuning.
+Bitune: Leveraging Bidirectional Attention to Improve Decoder-Only LLMs
+Dawid J.
+Kopiczkothanks: Correspondence: dawid.kopiczko@utn.de Fundamental AI Lab University of Technology Nuremberg          Tijmen Blankevoort Meta          Yuki M.
+Asano Fundamental AI Lab University of Technology Nuremberg
+Figure 1: Overview of Bitune.
+(a) During the prefilling phase, features are obtained from the prompt using both causal and bidirectional attention in two passes with separate weights.
+The two sets of keys and values are then combined using a weighted average before being passed to the decoding phase.
+(b) During the decoding phase, new tokens are generated in the standard way with causal attention, utilizing the features extracted from the instruction in the previous step, along with the features of other generated tokens.
 1 Introduction
-Large Language models (LLMs) are being used in many practical scenarios where humans interact with them through questions and answers.
-In applications such as general purpose assistants (OpenAI, 2024a), medical diagnosticians (Thirunavukarasu et al., 2023), game-conversation generation (Cox & Ooi, 2023) or coding-assistants (Roziere et al., 2023), the ability for an LLM to precisely follow an instruction and answer a question are of primary concern.
+Large Language Models (LLMs) are being deployed in numerous practical applications where humans engage with them through various forms of natural language interaction.
+In use cases such as general purpose assistants (OpenAI, 2024), medical diagnosticians (Thirunavukarasu et al., 2023), game-conversation generation (Cox and Ooi, 2023) or coding-assistants (Roziere et al., 2023), the ability for an LLM to precisely interpret and respond to user inputs is of primary concern.
 Correspondingly, Instruction-Tuning (IT)  (Chung et al., 2024; Ouyang et al., 2022a) is the prevailing paradigm for finetuning LLMs after their self-supervised pretraining phase to improve them for such tasks.
 Here, the model is trained on a dataset comprised of pairs of instructions and corresponding responses.
-Given the instruction-with-response structure of IT data, the generation of an LLM response can be divided into two phases: first, converting the instruction into Key and Value (KV) embeddings, stored in a KV-cache, which we refer to as instruction features; second, using these features to autoregressively generate an answer.
+Given the instruction-with-response structure of IT data, the generation of an LLM response can be divided into two phases: first, converting the instruction into key and value embeddings, which we refer to as instruction features; second, using these features to autoregressively generate an answer.
 Due to this task’s inherently conditional nature, the instruction features’ effectiveness is crucial for obtaining high-quality model outputs.
-In the past, bidirectional attention (Schuster & Paliwal, 1997) has been a key technique for obtaining stronger features for words or tokens.
+In the past, bidirectional attention (Schuster and Paliwal, 1997) has been a key technique for obtaining stronger features for words or tokens.
 This is because the meaning of a word depends greatly on its context.
 In particular, for some words in a sentence, the information that comes later might be far more informative for generating a meaningful representation and resolving ambiguities.
-But with only uni-directional causal attention, where the representation of each word is restricted to depend solely on the words that came before, this cannot be achieved.
-This is the reason why many previous transformers such as encoder-only BERT (Devlin et al., 2019) and encoder-decoder T5 (Raffel et al., 2020) employed bidirectional attention to improve the encoding of the input and why tasks like text retrieval (Lewis et al., 2020; Li & Li, 2023) and even the latest text-to-image generative models (OpenAI, 2024b; Esser et al., 2024) still rely on this.
-However, in the context of LLMs, architectures utilizing bidirectional attention have fallen out of favor, as decoder-only models such as GPT (OpenAI, 2024a) and Llama (AI@Meta, 2024) have focused on and vastly improved the generative performance of language models.
+With only uni-directional causal attention, where the representation of each word is restricted to depend solely on the words that came before, this cannot be achieved.
+This is the reason why many previous transformers such as encoder-only BERT (Devlin et al., 2019) and encoder-decoder T5 (Raffel et al., 2020) employed bidirectional attention to improve the encoding of the input and why tasks like text retrieval (Lewis et al., 2020; Li and Li, 2023) still rely on this.
+However, in the context of LLMs, architectures utilizing bidirectional attention have fallen out of favor, as decoder-only models such as GPT (OpenAI, 2024) and Llama (AI@Meta, 2024) have focused on and vastly improved the generative performance of language models.
 These architectures are trained by large volumes of data with next-token prediction, eschewing any look-ahead mechanism for the sake of better autoregressive modeling.
@@
 Overall, our contributions are as follows:
-  • We propose a novel method, Bitune, that improves the performance of pretrained decoder-only LLMs in question-answering scenarios.
+  • We propose a novel method, Bitune, that improves the performance of pretrained decoder-only LLMs in instruction-following and question-answering settings.
   • We evaluate the method on multiple downstream tasks, showing consistent improvements over the baselines.
-  • We conduct an extensive ablation study investigating the necessity of each component of the method, and showing the method’s PEFT-agnosticism.
-Figure 1: Overview of Bitune.
-(a) During the prefilling phase, features are obtained from the prompt using both causal and bidirectional attention in two passes with separate weights.
-The two sets of keys and values are then combined using a weighted average before being passed to the decoding phase.
-(b) During the decoding phase, new tokens are generated in the standard way with causal attention, utilizing the features extracted from the instruction in the previous step, along with the features of other generated tokens.
+  • We conduct an extensive ablation study investigating the necessity of each component of the method, and showing the method’s PEFT-agnosticism, as well as its effectiveness in full finetuning scenarios.
 2 Bidirectional Instruction-Tuning
-Algorithm 1 Python-like pseudocode of Bitune inference.
+Algorithm 1 Pseudocode of Bitune inference.
 # prompt - tensor with tokenized instruction
-# theta - tensor with mixing coeff. for each layer
-# theta_init - initial value of mixing coefficients
-k_c, v_c = model_causal(prompt) # Pass on the instruction for causal features
-k_b, v_b = model_bidir(prompt) # Pass to obtain bidirectional features
+# theta - tensor with mixing coefficients
+# theta_init - initial value of theta
+# Extract causal features
+k_c, v_c = model_causal(prompt)
+# Extract bidirectional features
+k_b, v_b = model_bidir(prompt)
 # Combine both sets of features
@@
 kv = (k, v)
-c_token = SEP # Initialize generation with a predefined token
+# Initialize generation with a predefined token
+c_token = SEP
 answer = [c_token]
-while c_token != EOS: # Stop generation on the end-of-sequence token
-# Get features of current token and logits of next token
+# Stop generation at the end-of-sequence token
+while c_token != EOS:
+# Get features of current token and logits
+# of next token
 k, v, logits = model_causal(c_token, kv)
-kv = concat(kv, (k, v)) # Concatenate obtained features with current KV cache
-c_token = get_token(logits) # Determine next predicted token based on logits
-answer.append(c_token) # Append generated token to the answer
+# Concatenate it with current KV cache
+kv = concat(kv, (k, v))
+# Determine next predicted token
+c_token = get_token(logits)
+# Append generated token to the answer
+answer.append(c_token)
 In the instruction-tuning setting (Ouyang et al., 2022b; Zhang et al., 2024), a dataset $\mathcal{D}$ consists of instruction-answer pairs that are used to adapt the model in a supervised fashion.
@@
 $p(a|q)=\Pi^{|a|}_{i=1}p(a_{i}|a_{1},\dots a_{i-1},q),$
-where $|a_{i}|$ denotes the length of the answer.
+where $|a|$ denotes the length of the answer.
 Note how compared to the regular language modeling objective, the response is already conditional (on the instruction $q$) even for the first generated token.
@@
 During the subsequent decoding phase, the model generates output tokens sequentially, one token at a time, based on the KV-cache of the instruction and the already generated tokens.
-In this work, we introduce, Bitune, a method to leverage this two-phase process to improve instruction-tuning of language models.
+In this work, we introduce Bitune, a method to leverage this two-phase process to improve instruction-tuning of language models.
 In our approach, the model processes the instruction with both causal and bidirectional attention using separate sets of parameters, leading to an enhanced KV-cache that is then used to condition the answer.
@@
 In other cases, they are the output of the preceding block and were processed by different components including the self-attention mechanism, which can be defined as:
-$\text{Attention}(Q,K,V,M)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V,$ (3)
-where $Q$ are the queries, $M$ is the attention mask, and $d_{k}$ is the dimension of keys and queries.
+$\text{Attn}(Q,K,V,M)=\sigma({QK^{T}}/{\sqrt{d_{k}}}+M)V,$ (3)
+where $\sigma$ is the softmax, $Q$ are the queries, $M$ is the attention mask, and $d_{k}$ is the dimension of keys and queries.
 For the causal pass, the mask $M_{\color[rgb]{1,0,0}c}$ enforces causality by masking future tokens, such that tokens $j$ can only attend to earlier tokens $i\leq j$, while for the bidirectional pass, no masking is applied:
-$M_{{{\color[rgb]{1,0,0}c}}}(i,j)=\begin{cases}0&\text{if }i\leq j\\ -\infty&\text{if }i>j\end{cases},\qquad M_{\color[rgb]{0,0,1}b}(i,j)=0.$ (4)
+$M_{{{\color[rgb]{1,0,0}c}}}(i,j)$ $=\begin{cases}0&\text{if }i\leq j\\ -\infty&\text{if }i>j\end{cases}$ (4)
+$M_{\color[rgb]{0,0,1}b}(i,j)$ $=0$ (5)
 The final KV-cache is obtained by a learnable convex combination of causal and bidirectional features,
-$K_{\texttt{Bitune}}=K_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{k})+K_{\color[rgb]{0,0,1}b}\cdot\alpha_{k},\quad V_{\texttt{Bitune}}=V_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{v})+V_{\color[rgb]{0,0,1}b}\cdot\alpha_{v},$ (5)
+$K_{\texttt{Bitune}}$ $=K_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{k})+K_{\color[rgb]{0,0,1}b}\cdot\alpha_{k}$ (6)
+$V_{\texttt{Bitune}}$ $=V_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{v})+V_{\color[rgb]{0,0,1}b}\cdot\alpha_{v},$ (7)
 where $\alpha$ represents the bidirectional-to-causal ratio of features. This ratio is parameterised as
-$\alpha_{j}=\frac{|\theta_{j}|}{\theta_{\text{init}}+|\theta_{j}|},\quad j\in\{k,v\}$ (6)
+$\alpha_{j}={|\theta_{j}|}/({\theta_{\text{init}}+|\theta_{j}|}),\quad j\in\{k,v\}$ (8)
 where $\theta_{j}$ is a learnable mixing coefficient per transformer block, and $\theta_{\text{init}}$ is a hyperparameter defining the initial value of $\theta_{j}$.
@@
 In the case of full finetuning, this approach would require an additional set of full weights, which is impractical for large models.
-Instead, we adapt our model to use bidirectional attention by using parameter-efficient finetuning methods.
+Instead, we adapt our model using parameter-efficient finetuning methods.
 These introduce only a fraction of trainable parameters, making it viable to have two modified variants of the model within a single forward pass.
@@
 (2022) to adapt the model.
-However, our method is PEFT-agnostic and can also utilize different methods for updating the weights, as demonstrated in our ablations section.
-In summary, Bitune is a method for improved instruction-tuning that learns a combination of causal and bidirectional features by applying two sets of lightweight PEFT methods to a pretrained model.
+However, Bitune can utilize different methods for updating the weights, including full model finetuning and other parameter-efficient techniques, as demonstrated in our ablations section.
 3 Experiments
-Table 1: Zero-shot results after instruction-tuning on the UltraFeedback dataset.
-We compare Bitune to the performance of the original model, the model finetuned with LoRA, and with LoRA16 using two times higher rank to match the parameter count of our method.
-Bitune significantly outperforms the baselines on almost all tasks for all models.
-PIQA ARC CSQA SIQA MMLU Avg.
-Model Method
-Gemma-2B Pretrained 57.5 36.9 35.5 38.2 34.0 40.4
-LoRA 66.7 43.4 42.3 44.3 31.7 45.7
-[…]
-Bitune 69.6 47.5 46.9 49.5 35.3 49.7
-Gemma-7B Pretrained 73.1 78.3 62.0 64.7 59.0 67.4
-LoRA 84.2 79.2 68.5 71.9 55.3 71.8
-[…]
-Bitune 83.6 80.1 69.2 72.7 53.8 71.9
-Llama2-7B Pretrained 59.2 38.1 32.6 45.1 36.0 42.2
-LoRA 69.5 49.9 45.3 57.0 41.1 52.6
-[…]
-Bitune 70.0 51.1 48.1 59.1 41.9 54.0
-Llama3-8B Pretrained 69.0 73.6 65.4 56.8 56.0 64.2
-LoRA 81.9 74.5 69.2 69.0 57.6 70.4
-[…]
-Bitune 84.4 77.4 72.7 70.1 59.0 72.7
-Phi-2 Pretrained 70.3 67.3 61.4 65.0 45.4 61.9
-LoRA 76.3 66.7 61.6 66.6 48.2 63.9
-[…]
-Bitune 76.5 67.2 63.0 68.5 48.9 64.8
 3.1 Instruction-Tuning
@@
 Specifically, we use a subset of the cleaned UltraFeedback (Cui et al., 2023) dataset, which contains instructions and corresponding answers generated by various LLMs.
-From this dataset, we select completions generated by GPT-4 (OpenAI, 2024a), ensuring high-quality responses for training.
-To fit every model on a single GPU, we filter out samples longer than 512 tokens, which leaves us with around 10,000 samples for training.
+From this dataset, we select completions generated by GPT-4 (OpenAI, 2024), ensuring high-quality responses for training.
+To fit every model on a single GPU, we filter out samples longer than 512 tokens, which leaves us with roughly 10,000 samples for training.
+For results on another instruction-tuning dataset, please see the Appendix A.6.
 We test the method on pretrained decoder-only language models of two different scales of approximately 2 billion and 7 billion parameters.
@@
 We use HuggingFace Transformers (Wolf et al., 2020) implementation of these models.
-We limit our experiments to these specific models due to their popularity and also our limited resources.
 […]
@@
 Note that this potentially puts our method at a disadvantage compared to the LoRA baseline.
-All hyperparameters are reported in the Appendix 6.2.
+All hyperparameters are reported in the Appendix A.2.
 Models are evaluated zero-shot on multiple-choice tasks to assess their performance.
@@
 As the tasks follow the question-answer pattern, they are compatible with the instruction-tuning setting.
+Table 1: Zero-shot results after instruction-tuning on the UltraFeedback dataset.
+We compare Bitune to the performance of the original model, the model finetuned with LoRA, and with LoRA16 using two times higher rank to match the parameter count of our method.
+Bitune significantly outperforms the baselines on almost all tasks for all models.
+Method PIQA ARC CSQA SIQA MMLU Avg.
+Gemma-2B
+Pretrained 57.5 36.9 35.5 38.2 34.0 40.4
+LoRA 66.7 43.4 42.3 44.3 31.7 45.7
+[…]
+Bitune 69.6 47.5 46.9 49.5 35.3 49.7
+Gemma-7B
+Pretrained 73.1 78.3 62.0 64.7 59.0 67.4
+LoRA 84.2 79.2 68.5 71.9 55.3 71.8
+[…]
+Bitune 83.6 80.1 69.2 72.7 53.8 71.9
+Llama2-7B
+Pretrained 59.2 38.1 32.6 45.1 36.0 42.2
+LoRA 69.5 49.9 45.3 57.0 41.1 52.6
+[…]
+Bitune 70.0 51.1 48.1 59.1 41.9 54.0
+Llama3-8B
+Pretrained 69.0 73.6 65.4 56.8 56.0 64.2
+LoRA 81.9 74.5 69.2 69.0 57.6 70.4
+[…]
+Bitune 84.4 77.4 72.7 70.1 59.0 72.7
+Phi-2
+Pretrained 70.3 67.3 61.4 65.0 45.4 61.9
+LoRA 76.3 66.7 61.6 66.6 48.2 63.9
+[…]
+Bitune 76.5 67.2 63.0 68.5 48.9 64.8
 For evaluation, we use the Language Model Evaluation Harness framework (Gao et al., 2023).
@@
 For our experiments, we opted to move the last token of the instruction template to the beginning of the modeled answer.
-For details on the instruction template used, please refer to the section 6.5 of the Appendix.
+For details on the instruction template used, please refer to the section A.10 of the Appendix.
 Results.
@@
 3.2 Downstream Task Training
-This complementary experiment verifies if Bitune increases the capacity of the model within the narrow scope of a single task.
+This complementary experiment verifies whether Bitune increases the capacity of the model within the narrow scope of a single task.
 It follows the setup from the instruction-tuning experiments with a few changes.
@@
 Results.
-Table 2: Result for the downstream task training.
+Table 2: Results for the downstream task training.
 We show accuracy on downstream tasks for the baseline LoRA finetuning and Bitune, averaged over 3 seeds.
@@
 We see the most significant gains on the GSM8K dataset, but slightly lower results for Gemma-7B and the SIQA task.
-PIQA ARC CSQA SIQA GSM8K Avg.
-Model Method
-Gemma-2B LoRA 81.4 58.0 77.2 77.4 30.2 64.8
+Method PIQA ARC CSQA SIQA GSM8K Avg.
+Gemma-2B
+LoRA 81.4 58.0 77.2 77.4 30.2 64.8
+[…]
 Bitune 83.3 60.0 78.3 76.6 33.0 66.2
-Gemma-7B LoRA 91.4 84.6 84.4 79.4 59.1 79.8
+Gemma-7B
+LoRA 91.4 84.6 84.4 79.4 59.1 79.8
+[…]
 Bitune 92.1 84.2 84.2 79.4 59.4 79.9
-Llama2-7B LoRA 84.4 66.6 81.5 82.7 32.0 69.4
+Llama2-7B
+LoRA 84.4 66.6 81.5 82.7 32.0 69.4
+[…]
 Bitune 84.4 66.9 82.0 81.4 32.9 69.5
-Llama3-8B LoRA 90.2 80.7 83.9 83.1 60.4 79.7
+Llama3-8B
+LoRA 90.2 80.7 83.9 83.1 60.4 79.7
+[…]
 Bitune 90.5 81.3 84.1 82.1 63.4 80.3
-Phi-2 LoRA 82.8 76.3 78.7 80.3 58.6 75.3
+Phi-2
+LoRA 82.8 76.3 78.7 80.3 58.6 75.3
+[…]
 Bitune 83.9 77.0 79.0 80.4 59.2 75.9
@@
 While there are a few cases where the baseline finetuning achieves better results on specific tasks, when considering the average gains, applying our method is beneficial across all models.
-Most importantly, on the GSM8K dataset, we see consistently high gains, suggesting that our method improves the model’s reasoning ability in generative tasks.
-The qualitative example for instruction-tuned model in Table 6 also supports this (more examples in Appendix 6.7).
+Most importantly, on the GSM8K dataset, we see consistent high gains, suggesting that our method improves the model’s reasoning ability in generative tasks.
+We present additional results on GSM8K with a 22B parameter model in the Appendix A.7.
 Similar to the instruction-tuning results, the highest gains are observed on the Gemma-2B model, while the lowest on the Gemma-7B.
 This indicates that the effectiveness of our method depends on the specific model used.
+3.3 Chain-of-Thought Reasoning
+To further test the effectiveness of our method on generative tasks that require explicit reasoning, we created four additional training sets for PIQA, ARC, CSQA, and GSM8K.
+Each set was distilled from the GPT-4.1 model (OpenAI, 2025) by prompting it to solve every training-split question step-by-step; the resulting reasoning traces were then used to finetune Gemma-2B and Llama-3-8B on each benchmark separately.
+At evaluation time the models generated a full answer, including intermediate reasoning, and scores were averaged over three random seeds.
+The original versions of these benchmarks supply only the final label, forcing a model trained on them to perform all reasoning in latent space and yielding short generations that do not reflect real-world usage.
+Our distilled chain-of-thought data instead mirrors the way large language models are typically invoked, producing longer, multi-step explanations before the answer, thereby providing a more faithful test bed.
+Moreover, because the public test sets may already appear in pre-training corpora, training the models to answer the same questions in a custom, step-by-step format reduces the risk that apparent gains stem from data contamination.
+The prompts used to distill the datasets, along with example completions, are available in Appendix A.9.
+Results.
+Table 3 shows that Bitune improves performance on chain-of-thought reasoning tasks, raising average accuracy by 2.4 pp for Gemma-2B and 1.3 pp for Llama-3-8B.
+Table 3: Results for the chain-of-thought training.
+Models were trained on downstream tasks with chain-of-thought reasoning traces distilled from GPT-4.1 model, and evaluated by generating the reasoning steps before providing the final answer to given question.
+Across the tasks and models tested, Bitune leads to significant gains when compared to LoRA baselines.
+Method PIQA ARC CSQA GSM8K Avg.
+Gemma-2B
+LoRA 77.5 55.4 63.2 40.3 59.1
+[…]
+Bitune 78.4 56.6 67.1 43.7 61.5
+Llama3-8B
+LoRA 88.2 82.5 78.6 70.3 79.9
+[…]
+Bitune 89.6 82.8 79.0 73.5 81.2
+3.4 Inference Speed Comparison
+During inference, Bitune performs two forward passes over the instruction sequence and merges the resulting features, which affects prefilling runtime.
+We compare inference time for Gemma-2B on a long-context instruction of 2000 tokens and report the time required to prefill this instruction and to generate 2000 subsequent tokens on a single A100 GPU.
+Table 4: Inference-time comparison for Gemma-2B. We report the time required to prefill a 2000-token instruction and to generate a further 2000 tokens on a single A100 GPU.
+Method Prefill Generation
+LoRA 0.03s 62.15s
+Bitune 0.28s 62.28s
+Table 4 shows that the extra latency for processing the instruction is negligible because most of the computation time is spent on autoregressive answer generation.
 3.5 Ablations
@@
 We see that all ablated variants outperform the LoRA baseline, and combining all components performs the best.
-Of note, especially bidirectional attention improves results the most.
-Avg. Acc.
-Model Method Causal Bidir. Mixing Sep. Weights
-Gemma-2B LoRA - - - 45.7
+Method Causal Bidir. Mixing Sep. weights Acc.
+Gemma-2B
+LoRA - - - 45.7
 Naive Bidir. - - - 47.9
@@
 Bitune 49.7
-Llama3-8B LoRA - - - 70.4
+Llama3-8B
+LoRA - - - 70.4
 Naive Bidir. - - - 71.9
@@
 - To answer this question, we do not introduce the second set of weights and use the same LoRA for both passes on the prompt, calling it Shared Weights.
-The results, averaged over three seeds and presented in Table 3, indicate that all variants of the method lead to gains over the baseline LoRA finetuning.
+The results, averaged over three seeds and presented in Table 5, indicate that all variants of Bitune lead to gains over the baseline LoRA finetuning.
 However, the highest gains are observed in the full variant of Bitune, demonstrating that each component contributes to the method’s effectiveness.
@@
 Table 6: Combining Bitune with different PEFT methods. Performances are averaged over 3 seeds. We can see that our method improves results regardless of the specific PEFT method used.
-PIQA ARC CSQA SIQA MMLU Avg.
-Model Method
-Gemma-2B LoRA 66.7 43.4 42.3 44.3 31.7 45.7
+Method PIQA ARC CSQA SIQA MMLU Avg.
+Gemma-2B
+LoRA 66.7 43.4 42.3 44.3 31.7 45.7
 Bitune 69.6 47.5 46.9 49.5 35.3 49.7
@@
 […]
-Llama3-8B LoRA 81.9 74.5 69.2 69.0 57.6 70.4
+Llama3-8B
+LoRA 81.9 74.5 69.2 69.0 57.6 70.4
 Bitune 84.4 77.4 72.7 70.1 59.0 72.7
@@
 The results are shown in Table 6.
-We find consistent gains across all three PEFT methods we analyze, with gains ranging from $+1.6\%$ to $+4.0\%$ for averaged accuracies.
+We find consistent gains across all three PEFT methods we analyze, with gains ranging from $1.6\%$ to $4.0\%$ for averaged accuracy.
 This demonstrates that Bitune is PEFT-agnostic and can be combined with existing and future innovations in PEFT methods.
-Initialization of Mixing Coefficient
-The initial value of the mixing coefficient $\theta$ is a hyperparameter in our method.
-To evaluate its impact on the performance and the training dynamics of the bidirectional-to-causal ratio of features, we conduct experiments on the instruction-tuning setup with the following values: $0.1$, $0.01$, and $0.001$.
-Figure 2: Bidirectional-to-causal ratio during training.
-The ratio is averaged over all layers and shown for different initial values of mixing coefficients for Llama3-8B.
-The initial value impacts the change of the ratio, with higher values slowing it down, and lower values increasing it.
-Figure 3: Ratio across layers. Here we show the final ratio of the model in Fig. 2 across all layers for the K and V values. The utilization of bidirectional attention is spread across all layers.
-Table 5(a) demonstrates that the initial value of the mixing coefficient impacts the performance, with $0.01$ being the most optimal value for both models, regardless of their scale.
-Figure 2 shows that the initial value substantially affects the rate of change of the mixing ratio, with the higher value leading to nearly no change in the ratio, while the lower value results in sharp changes at the very beginning of the training.
-From Figure 3, we also observe that after training, all layers utilize the bidirectional attention.
+Full Finetuning
+Additionally, we test whether Bitune leads to gains with full finetuning (Full-FT), by optimizing two sets of full model’s parameters.
+We conduct experiments on Gemma-2B model, and compare results with standard Full-FT baseline.
+The results in Table 7 demonstrate that Bitune improves the model’s performance even in full finetuning scenarios.
+Table 7: Results for full finetuning Gemma-2B model on the instruction-tuning setup. We compare our method to full finetuning baseline. For Bitune, we optimize two sets of full model’s weights.
+Method PIQA ARC CSQA SIQA MMLU Avg.
+Full FT 69.0 46.7 43.5 43.7 34.8 47.5
+[…]
 Attention Mask of Second Pass
@@
 We transpose the causal attention mask, blocking information flow from the past tokens, and allowing from the future tokens - we call it anti-causal attention mask.
-Table 5: Ablation of Bitune’s attention.
-We vary the additional attention mask for processing the instruction besides the causal pass, and evaluate different initial values of mixing coefficient, $\theta_{\text{init}}$.
-(a) Different init. values for mixing coefficients.
-Model Init. Value Avg. Acc.
-Gemma-2B $0.1$ 49.4
-$0.01$ 49.7
-$0.001$ 47.2
-Llama3-8B $0.1$ 72.7
-$0.01$ 72.7
-$0.001$ 72.3
-(b) 2nd pass attention masks for instruction features.
+Table 8: Ablation of Bitune’s attention. We vary the attention mask for extracting the second set of features.
 Model Attention Mask Avg. Acc
@@
 Bidirectional 72.7
-Results shown in Table 5(b) indicate that the instruction has to be processed with full bidirectional attention to achieve the highest gains.
+Results shown in Table 8 indicate that the instruction has to be processed with full bidirectional attention to achieve the highest gains.
 Combining causal and anti-causal features independently does not lead to the same high performance.
-Table 6: Qualitative results on Llama3-8B. We show a response for LoRA and Bitune.
-[GSM8K]     Question: Janet’s ducks lay 16 eggs per day.
-She eats three for breakfast every morning and bakes muffins for her friends every day with four.
-She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg.
-How much in dollars does she make every day at the farmers’ market?
-LoRA Janet makes \$16 - (3 + 4) = \$9 per day at the farmers’ market.
-Bitune Janet makes \$2 per fresh duck egg, and she sells the remainder of the eggs she collects each day.
-If she eats three eggs for breakfast and bakes muffins with four eggs, she has 16 - 3 - 4 = 9 eggs left to sell.
-So, she makes \$2 x 9 = \$18 every day at the farmers’ market.
 4 Related Work
@@
 Additionally, UL2 (Tay et al., 2023) introduces a pretraining objective called "Mixture of Denoisers", which combines various denoising strategies, including the prefix-LM approach.
-Lastly, XLNet Yang et al.
-(2019) also allows for non-causal word ordering by allowing random permutations to be used with a next-token prediction objective.
+Lastly, XLNet (Yang et al., 2019) also allows for non-causal word ordering by allowing random permutations to be used with a next-token prediction objective.
 All these works focused on the model pretraining.
@@
 It can be done by enabling bidirectional attention, training the model on the objective of masked token prediction, and applying unsupervised contrastive learning.
-5 Discussion
-Limitations.
+Bitune vs. Prefix-LM.
+Both Bitune and the classic prefix-LM masking scheme grant full bidirectional access to the instruction prefix.
+The key difference is that Bitune is a hybrid approach designed for the optimal adaptation of decoder-only models.
+Existing large language models are almost always pre-trained as pure decoder-only networks; simply replacing the causal mask with the prefix-LM masking scheme before finetuning therefore discards the causal features encoded during pretraining and forces the model to relearn the mechanics of bidirectional attention over the prefix tokens.
+Bitune keeps the original causal stream intact and adds a second bidirectional pass, injecting fresh capacity into the model.
+Finetuning then proceeds in a single architecture that unifies the two attention regimes.
+Our ablation study (Section 3.5) confirms the payoff: Bitune yields larger gains than straightforward “prefix-LM conversion” (ablation Naive Bidir).
+Whether the same advantage holds when both schemes are trained from scratch remains an open question.
+5 Conclusion
+This work proposes a method that exploits the inherent instruction–answer structure of IT datasets to incorporate bidirectional attention into pretrained decoder-only models.
+Bitune demonstrates general applicability across different models and scales, and it consistently improves performance across commonsense reasoning, arithmetic, and language understanding tasks.
+We further demonstrate that the method is compatible with different existing PEFT methods and will likely benefit from further developments in this area.
+Limitations
 During standard instruction-tuning training, the instruction and the answer are processed in a single forward pass.
-In our method, this processing is explicitly split into phases, extracting instruction features with two passes & answer modeling, increasing both training time and memory usage.
-As for the inference, these two passes on the instruction, bidirectional and causal, can be processed either in parallel, impacting memory usage, or sequentially, increasing latency.
+In our method, this processing is explicitly split into phases - extracting instruction features with causal attention, bidirectional attention, and answer modeling, which results in increased both training time and memory usage.
 However, this is a minor limitation in the context of instruction-tuning, since typically smaller datasets are used compared to pretraining, leading to relatively short training times.
-Furthermore, at inference time the added latency for processing the instruction is negligible, as the bulk of compute is used for the autoregressive answer generation: e.g.
-only a 0.2s increase from 11.5s to 11.7s for a 200 token response to a 50 token instruction for Llama3-8B (see Appendix 6.4 for more results).
-Conclusion.
-Bitune is the first method to utilize the instruction-answer structure of IT datasets to propose a new finetuning method that enables bidirectional processing in pretrained decoder-only models.
-Bitune demonstrates general applicability across different models and scales.
-It consistently improves performance across commonsense reasoning, arithmetic, and language understanding tasks.
-We further demonstrate that our method is compatible with different existing PEFT methods and will likely benefit from further developments in this era.
-As the first work to demonstrate the potential of an instruction-tuning-specific adaptation for decoder LLMs, we believe there is much space for further research into this promising and crucial direction.
-Acknowledgements
-We acknowledge the use of the National Supercomputer Snellius and the Distributed ASCI Supercomputer 6 (Bal et al., 2016) for essential computational tasks.
+Furthermore, as shown in Section 3.4, the additional latency during prefilling at inference time is negligible, since most time is spent on sequential token generation.
+Appendix A Appendix
+A.1 Pseudocode for Bitune Training Step
+Algorithm 2 Python-like pseudocode of Bitune Training Step.
+# prompt - tensor with tokenized instruction
+# answer - tensor with tokenizer answer to model
+# theta - trainable tensor with mixing coeff. for each layer
+# theta_init - initial value of mixing coefficients
+k_c, v_c = model_causal(prompt) # Pass on the instruction for causal features
+k_b, v_b = model_bidir(prompt) # Pass to obtain bidirectional features
+# Combine both sets of features
+alpha = theta.abs() / (theta.abs() + theta_init)
+k = k_c * (1 - alpha) + k_b * alpha
+v = v_c * (1 - alpha) + v_b * alpha
+kv = (k, v)
+logits = model_causal(answer, kv)
+loss = compute_loss(logits, answer)
+loss.backward()
+update_parameters(model_causal, model_bidir, theta)
+A.2 Hyperparameters
+Table 9: Hyperparameters shared across models and datasets.
+Hyperparameter Value
+GPUs 1
+Optimizer AdamW
+LR Scheduler Linear
+Weight Decay 0.0
+Batch Size (incl. accumulation) 10
+Accumulation Steps 10
+Warmup Steps 10% of update steps
+$\theta_{\text{init}}$ (Bitune) 0.01
+RNG Seeds 42, 43, 44
+Table 10: Dataset-specific hyperparameters.
+Hyperparameter UltraFeedback PIQA ARC CSQA SIQA GSM8K
+Epochs 3 1 5 1 1 1
+Update Steps 3000 1605 555 974 3341 747
+Table 11: Learning rate for given dataset-model pair, including different PEFT variants for instruction-tuning experiments.
+Model UltraFeedback PIQA ARC CSQA SIQA GSM8K
+Gemma-2B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
+Gemma-7B 3E-4 3E-4 3E-4 3E-4 3E-4 3E-4
+Llama2-7B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
+Llama3-8B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
+Phi-2 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
+Gemma-2B (DoRA) 3E-4 - - - - -
+Gemma-2B (IA3) 1E-3 - - - - -
+Llama3-8B (DoRA) 3E-4 - - - - -
+Llama3-8B (IA3) 1E-3 - - - - -
+Table 12: Configuration of PEFT methods. All other hyperparameters have default values of HuggingFace PEFT library.
+Hyperparameter Value
+Rank (LoRA, DoRA) 8
+Alpha (LoRA, DoRA) 1
+Target Modules (All) All linear layers of MLP and Self-Attention
+Feedforward Modules (IA3) All linear layers of MLP
+A.3 Datasets
+For all experiments we used HuggingFace Datasets (Lhoest et al., 2021) library to obtain necessary datasets.
+Table 13: Table with datasets and corresponding paths, to be used with HuggingFace Datasets library.
+Dataset Path
+UltraFeedback openbmb/UltraFeedback
+PIQA piqa
+ARC-Challenge allenai/ai2_arc
+CSQA tau/commonsense_qa
+SIQA social_i_qa
+GSM8K gsm8k
+MMLU hails/mmlu_no_train
+A.4 Training Speed & Memory Usage
+As the method introduces two additional forward passes during training, both the training speed and the memory usage are impacted.
+Here we present average training times and GPU memory usage on the instruction-tuning setup with 3000 update steps (30000 actual steps, due to gradient accumulation), on a single A100 GPU, for models of two different scales - Gemma-2B & Llama3-8B.
+Our implementation has not been optimized, which means that e.g.
+training times could be improved via parallelization of two passes on the prompt.
+Table shows average training time, peak GPU memory usage during training, and average accuracy on downstream tasks.
+Train. Time [h] Memory [GB] Acc.
+Model Method
+Gemma-2B LoRA 1.0 14.9 45.7
+Bitune 3.1 19.8 49.7
+Llama3-8B LoRA 1.7 26.6 70.4
+Bitune 5.3 30.8 72.7
+Using these values, one can approximate required compute to reproduce results on a given tasks, as all experiments shared the same batch size and many other hyperparameters.
+A.5 Initialization of Mixing Coefficient
+The initial value of the mixing coefficient $\theta$ is a hyperparameter in our method.
+To evaluate its impact on the performance and the training dynamics of the bidirectional-to-causal ratio of features, we conduct experiments on the instruction-tuning setup with the following values: $0.1$, $0.01$, and $0.001$.
+Figure 2: Bidirectional-to-causal ratio during training.
+The ratio is averaged over all layers and shown for different initial values of mixing coefficients for Llama3-8B.
+The initial value impacts the change of the ratio, with higher values slowing it down, and lower values increasing it.
+Figure 3: Ratio across layers. Here we show the final ratio of the model in Fig. 2 across all layers for the K and V values. The utilization of bidirectional attention is spread across all layers.
+Model Init. Value Avg. Acc.
+Gemma-2B $0.1$ 49.4
+$0.01$ 49.7
+$0.001$ 47.2
+Llama3-8B $0.1$ 72.7
+$0.01$ 72.7
+$0.001$ 72.3
+Table 15: Different init. values for mixing coefficients.
+Table 15 demonstrates that the initial value of the mixing coefficient impacts the performance, with $0.01$ being the most optimal value for both models, regardless of their scale.
+Figure 2 shows that the initial value substantially affects the rate of change of the mixing ratio, with the higher value leading to nearly no change in the ratio, while the lower value results in sharp changes at the very beginning of the training.
+In Figure 3, we also observe that after training, all layers utilize the bidirectional attention.
+A.6 Instruction-tuning on Alpaca Dataset
+We tested Bitune on another, larger instruction-tuning dataset - cleaned Alpaca dataset [^1] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023) with 50,000 samples.
+Similarly to previous experimental settings, first we tested different learning rates for the baseline LoRA finetuning, picked the best one, and then used the same learning rate for other approaches used in the experiment - Bitune, and Naive Bidir.
+(introduced in the ablation study section 3.5).
+The results demonstrate that our method’s benefits extend to larger datasets as well.
+Table 16: Zero-shot results after instruction-tuning on the Alpaca dataset. Results are averaged over 3 random seeds.
+PIQA ARC CSQA SIQA MMLU Avg.
+Model Method
+Gemma-2B LoRA 64.1 40.8 38.6 38.6 33.8 43.2
+Naive Bidir. 64.9 41.7 42.2 42.3 33.6 44.9
+Bitune 66.5 43.6 43.2 43.5 36.2 46.6
+Llama3-8B LoRA 76.6 71.1 63.3 67.4 57.8 67.2
+Naive Bidir. 79.5 70.5 62.4 66.8 55.3 66.9
+Bitune 79.4 71.3 63.5 68.6 58.3 68.2
+A.7 Finetuning Larger Model on GSM8K
+In order to verify whether improvements hold for larger, already highly capable models, we finetune Codestral (Mistral AI, 2024) with 22B parameters on the GSM8K dataset.
+The results, averaged over 3 random seeds, show a substantial 4.3 percentage point improvement over the baseline LoRA finetuning, indicating that Bitune can provide considerable gains even for larger models, which already have strong performance on a given task.
+Table 17: Results for Bitune and baseline LoRA finetuning on GSM8K dataset.
+GSM8K
+Model Method
+Codestral-22B LoRA 69.3
+Bitune 73.6
+A.8 Multi-Turn Setting
+In our experiments, we focus on single-turn QA settings, but the approach can be extended to multi-turn chat applications in the following two ways:
+(A) “Recompute”  The most straightforward way is to treat the entire chat history as a prefix or query and recompute the KV cache at every new round.
+This allows bidirectional attention over all tokens, including previously generated answers.
+(B) “Alternating”  Another option is to alternate attention patterns with the following procedure:
+  1. 1.
+    Prefill the first instruction and generate the initial KV cache by merging causal and bidirectional features.
+  2. 2.
+    Generate output token by token, appending the KV of each output to the cache (the output KV uses only causal features).
+  3. 3.
+    Prefill the next instruction with both attention masks and append its KV to the cache.
+  4. 4.
+    Repeat from step 2.
+A.9 Chain of Thought Data Generation
+For each train set of datasets - PIQA, ARC-Challenge, CSQA, GSM8K - we have distilled chain-of-thought completions with GPT-4.1 model.
+The completions were asked to be in a specific format allowing to extract the steps, and final answer separately. We have used the following prompt templates:
+Dataset Prompt
+PIQA Think step-by-step, then provide the final answer.
+Separate each operation / thought into a single statement.
+Try to make every statement and whole process as concise as possible.
+Use only a single letter for the final answer.
+If multiple answer seem likely to be correct, pick the most likely one.
+Use the following structure (with example statements): <step> first reasoning step goes here </step> <step> second reasoning step goes here </step> <step> and so on… </step> <answer> B </answer> [query]
+ARC Think step-by-step, then provide the final answer.
+Separate each operation / thought into a single statement.
+Try to make every statement and whole process as concise as possible.
+Use only a single letter for the final answer.
+If multiple answer seem likely to be correct, pick the most likely one.
+Use the following structure (with example statements): <step> first reasoning step goes here </step> <step> second reasoning step goes here </step> <step> and so on… </step> <answer> C </answer> [query]
+CSQA Think step-by-step, then provide the final answer.
+Separate each operation / thought into a single statement.
+Try to make every statement and whole process as concise as possible.
+Use only a single letter for the final answer.
+If multiple answer seem likely to be correct, pick the most likely one.
+Use the following structure (with example statements): <step> Identify the nature of a drawstring bag, which is typically small and used for carrying personal or valuable items.
+</step> <step> Consider the context of checking baggage, which is commonly done at airports.
+</step> <step> Eliminate options that are less likely related to checking baggage: garbage can and safe (unlikely travel destinations).
+</step> <step> Assess the options related to travel or specific locations: military, jewelry store, airport.
+</step> <step> Military is possible, but typically does not involve checking baggage as a drawstring bag.
+</step> <step> Jewelry store is a destination but not typically associated with checking baggage during travel.
+</step> <step> Airport is the most logical location where baggage is checked, including a drawstring bag.
+</step> <answer> E </answer> [query]
+GSM8K Think step-by-step, then provide the final answer.
+Separate each operation and calculation into one statement.
+Try to make every statement and whole process as concise as possible.
+For the final answer, provide the integer value only, without any units.
+Use the following structure: <step> first reasoning step goes here… </step> <step> second reasoning step goes here… </step> <step> and so on… </step> <answer> 42 </answer> [query]
+Here we provide the example completions:
+Dataset Example Completion
+PIQA Question: When boiling butter, when it’s ready, you can Choices: A.
+Pour it onto a plate B.
+Pour it into a jar Steps: Boiling butter usually refers to melting and clarifying it by heating.
+When clarified butter is ready, it is typically poured into a container to cool and solidify.
+A jar is more appropriate for storing melted or clarified butter.
+Pouring onto a plate is uncommon and impractical for storage or further use.
+Answer: B
+ARC Question: George wants to warm his hands quickly by rubbing them.
+Which skin surface will produce the most heat?
+Choices: A: dry palms B: wet palms C: palms covered with oil D: palms covered with lotion Steps: Friction generates heat when two surfaces rub against each other.
+Dry skin produces more friction compared to wet or oily surfaces because water or oil acts as a lubricant.
+Oil and lotion reduce friction by creating a slippery layer, so less heat is produced.
+Wet palms also reduce friction compared to dry palms because water acts as a lubricant.
+Therefore, dry palms will produce the most heat when rubbed together.
+Answer: A
+CSQA Question: The sanctions against the school were a punishing blow, and they seemed to what the efforts the school had made to change?
+Choices: A: ignore B: enforce C: authoritarian D: yell at E: avoid Steps: Understand the sentence structure: ’The sanctions … seemed to … the efforts the school had made to change.’ The word to fill should describe what the sanctions did to the school’s efforts.
+Sanctions usually negatively impact or counteract efforts rather than support them.
+Look at the choices: ignore, enforce, authoritarian, yell at, avoid.
+Eliminate irrelevant options: ’authoritarian’ is an adjective, ’yell at’ is informal and unsuitable, ’enforce’ contradicts the negative impact idea.
+Consider ’ignore’ and ’avoid’ both suggest neglect or dismissal, which fits the idea of a punishing blow counteracting efforts.
+’Ignore’ means to pay no attention; ’avoid’ means to keep away from actively.
+’Ignore’ better fits the idea of sanctions negating or discounting the efforts, making it the most likely choice.
+Answer: A
+GSM8K Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May.
+How many clips did Natalia sell altogether in April and May?
+Steps: Identify the number of clips sold in April: 48.
+Calculate the number of clips sold in May as half of April’s sales: 48 / 2 = 24.
+Calculate the total number of clips sold in April and May by adding both amounts: 48 + 24 = 72.
+Answer: 72
+A.10 Prompt Templates
+Templates used to format instruction-answer pairs for a given dataset, for both training and evaluation. In all cases there is a space character at the beginning of the answer part.
+Dataset Instruction Answer
+UltraFeedback Question: {instruction}
+Answer:
+{completion}<EOS>
+PIQA Question: {question}
+Choices:
+{choice0}
+{choice1}
+Answer:
+{answer}<EOS>
+ARC Question: {question}
+Choices:
+{choice0}
+{choice1}
+{choice2}
+{choice3}
+Answer:
+{answer}<EOS>
+CSQA Question: {question}
+Choices:
+{choice0}
+{choice1}
+{choice2}
+{choice3}
+{choice4}
+Answer:
+{answer}<EOS>
+SIQA Question: Given the context, answer correctly the question.
+Context: {context}
+Question: {question}
+Choices:
+(0) {choice0}
+(1) {choice1}
+(2) {choice2}
+Answer:
+({answer_index})<EOS>
+GSM8K Question: {question}
+Answer:
+{answer}<EOS>
+MMLU Question: {question}
+(0). {choice0}
+(1). {choice1}
+(2). {choice2}
+(3). {choice3}
+Answer:
+({answer_index})<EOS>
+A.11 Results with Standard Deviation
+Tables with complete results averaged over 3 seeds, includes standard deviation.
+Table 21: Zero-shot results on downstream tasks after instruction-tuning on the UltraFeedback dataset.
+PIQA ARC CSQA SIQA MMLU
+mean std mean std mean std mean std mean std
+Model Method
+Gemma-2B Pretrained 57.51 - 36.86 - 35.46 - 38.18 - 34.01 -
+LoRA 66.74 0.52 43.37 1.33 42.32 1.60 44.27 3.85 31.74 0.78
+[…]
+Bitune 69.59 1.20 47.47 0.64 46.87 1.98 49.51 1.54 35.29 0.08
+Gemma-7B Pretrained 73.12 - 78.33 - 62.00 - 64.74 - 59.04 -
+LoRA 84.24 0.51 79.18 0.76 68.47 1.61 71.92 0.34 55.26 0.43
+[…]
+Bitune 83.59 0.46 80.09 0.90 69.15 0.56 72.74 0.90 53.81 0.40
+Llama2-7B Pretrained 59.25 - 38.14 - 32.60 - 45.09 - 35.98 -
+LoRA 69.51 0.73 49.94 0.79 45.32 2.65 57.05 1.24 41.06 0.23
+[…]
+Bitune 70.00 0.53 51.11 0.23 48.08 2.59 59.09 0.96 41.87 0.21
+Llama3-8B Pretrained 68.99 - 73.63 - 65.36 - 56.81 - 56.00 -
+LoRA 81.94 0.38 74.46 1.06 69.23 0.67 68.99 1.55 57.62 0.54
+[…]
+Bitune 84.39 0.24 77.42 1.15 72.70 0.82 70.15 0.34 58.96 0.37
+Phi-2 Pretrained 70.35 - 67.32 - 61.43 - 65.05 - 45.41 -
+LoRA 76.31 0.17 66.67 0.44 61.62 0.40 66.65 0.31 48.23 0.75
+[…]
+Bitune 76.51 0.30 67.18 0.49 63.04 0.71 68.51 0.56 48.92 0.41
+Table 22: Result for the downstream task training setup.
+PIQA ARC CSQA SIQA GSM8K
+mean std mean std mean std mean std mean std
+Model Method
+Gemma-2B LoRA 81.41 0.46 58.05 1.16 77.18 0.05 77.40 0.40 30.17 0.75
+Bitune 83.28 0.33 59.98 1.19 78.32 0.68 76.60 0.24 32.98 0.97
+Phi-2 LoRA 82.79 1.15 76.31 0.56 78.68 0.76 80.30 0.44 58.55 0.89
+Bitune 83.91 0.28 77.02 0.34 78.95 0.22 80.38 0.21 59.21 1.15
+Gemma-7B LoRA 91.42 0.60 84.58 0.36 84.41 1.04 79.36 1.41 59.11 0.97
+Bitune 92.13 0.27 84.22 0.74 84.22 0.76 79.39 1.33 59.39 0.72
+Llama3-8B LoRA 90.24 0.27 80.75 0.36 83.92 0.74 83.15 0.41 60.45 0.72
+Bitune 90.52 0.78 81.26 0.79 84.11 1.05 82.12 0.30 63.43 0.27
+Llama2-7B LoRA 84.39 0.69 66.55 0.78 81.52 0.83 82.67 0.31 32.02 0.68
+Bitune 84.39 0.61 66.87 0.90 81.95 0.56 81.39 0.15 32.85 1.05
+Table 23: Ablation study on the components of Bitune.
+PIQA ARC CSQA SIQA MMLU
+mean std mean std mean std mean std mean std
+Model Method
+Gemma-2B LoRA 66.74 0.52 43.37 1.33 42.32 1.60 44.27 3.85 31.74 0.78
+Bitune 69.59 1.20 47.47 0.64 46.87 1.98 49.51 1.54 35.29 0.08
+Naive Bidir. 67.79 1.00 44.65 1.78 46.79 2.86 48.04 0.95 32.43 1.40
+No Mixing 69.01 1.47 45.71 1.32 46.14 2.56 49.80 0.53 34.03 0.35
+Only Causal 66.39 1.28 45.28 0.65 42.45 1.65 46.98 0.76 33.27 0.41
+Shared Weights 68.10 0.41 44.34 0.90 44.53 1.67 47.19 2.16 32.93 0.13
+Llama3-8B LoRA 81.94 0.38 74.46 1.06 69.23 0.67 68.99 1.55 57.62 0.54
+Bitune 84.39 0.24 77.42 1.15 72.70 0.82 70.15 0.34 58.96 0.37
+Naive Bidir. 85.44 0.25 76.45 0.60 69.37 0.28 70.04 0.43 58.33 0.89
+No Mixing 85.56 0.38 74.86 0.21 68.60 1.31 69.74 0.11 58.88 0.43
+Only Causal 82.37 0.41 74.63 0.91 70.65 0.93 69.17 1.51 58.43 0.26
+Shared Weights 84.10 0.63 75.94 1.11 71.91 0.71 70.61 0.60 59.03 0.79
+Table 24: Results for different PEFT methods used in combination with Bitune.
+PIQA ARC CSQA SIQA MMLU
+mean std mean std mean std mean std mean std
+Model Method
+Gemma-2B LoRA 66.74 0.52 43.37 1.33 42.32 1.60 44.27 3.85 31.74 0.78
+Bitune 69.59 1.20 47.47 0.64 46.87 1.98 49.51 1.54 35.29 0.08
+DoRA 66.70 0.63 43.57 0.91 41.88 1.08 44.71 3.83 31.95 0.72
+[…]
+IA3 67.25 0.52 46.50 0.67 45.54 0.30 37.65 1.07 32.49 0.20
+[…]
+Llama3-8B LoRA 81.94 0.38 74.46 1.06 69.23 0.67 68.99 1.55 57.62 0.54
+Bitune 84.39 0.24 77.42 1.15 72.70 0.82 70.15 0.34 58.96 0.37
+DoRA 82.12 0.44 75.40 0.52 70.22 0.39 69.19 1.14 57.67 0.23
+[…]
+IA3 80.87 0.58 75.54 0.10 68.25 0.26 66.38 0.18 58.75 0.24
+[…]
+Table 25: Results for different init values for mixing coefficients.
+PIQA ARC CSQA SIQA MMLU
+mean std mean std mean std mean std mean std
+Model Init. Value
+Gemma-2B 0.1 69.68 0.74 46.33 0.81 45.97 2.64 49.86 2.21 35.25 0.46
+0.01 69.59 1.20 47.47 0.64 46.87 1.98 49.51 1.54 35.29 0.08
+0.001 67.79 0.88 44.20 0.68 45.56 1.61 46.35 1.08 31.89 0.67
+Llama3-8B 0.1 85.00 0.73 76.48 0.63 71.06 1.56 71.68 0.46 59.19 0.38
+0.01 84.39 0.24 77.42 1.15 72.70 0.82 70.15 0.34 58.96 0.37
+0.001 84.60 0.62 76.62 0.09 71.28 1.45 70.62 0.15 58.49 0.22
+A.12 Example Attention Matrices
+Here we show pairs of matrices with causal and bidirectional attention scores of Bituned Gemma-2B for the first two GSM8K samples, with scores averaged over layers and heads.
+Source tokens are represented with the vertical axis, while target tokens (for which the attention is paid to) with the horizontal one - e.g.
+the first column represents attention paid to the first (BOS) token, by each other token.
+Darker color represents higher attention score.
+Causal Bidirectional
+[…]
+[Uncaptioned image] [Uncaptioned image]
+[…]
+[Uncaptioned image] [Uncaptioned image]
+A.13 GSM8K Samples
+Samples of responses to the first 20 questions from GSM8K benchmark. Shown for Llama3-8B intruction-tuned on UltraFeedback dataset - for standard finetuning with LoRA and Bitune.
+Prompt LoRA Bitune
+Question: Janet’s ducks lay 16 eggs per day.
+She eats three for breakfast every morning and bakes muffins for her friends every day with four.
+She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg.
+How much in dollars does she make every day at the farmers’ market?
+Janet makes \$16 - (3 + 4) = \$9 per day at the farmers’ market.
+Janet makes \$2 per fresh duck egg, and she sells the remainder of the eggs she collects each day.
+If she eats three eggs for breakfast and bakes muffins with four eggs, she has 16 - 3 - 4 = 9 eggs left to sell.
+So, she makes \$2 x 9 = \$18 every day at the farmers’ market.
+Question: A robe takes 2 bolts of blue fiber and half that much white fiber.
+How many bolts in total does it take?
+A robe requires 2 bolts of blue fiber and half that much white fiber, which is equal to 1 bolt of white fiber.
+Therefore, the total number of bolts required for a robe is 3 bolts (2 blue and 1 white).
+To make a robe, you need 2 bolts of blue fiber and half that much white fiber.
+So, the total number of bolts required is: 2 bolts of blue fiber + half of 2 bolts of blue fiber (which is 1 bolt of white fiber) Total bolts = 2 + 1 = 3 bolts So, the total number of bolts required to make a robe is 3 bolts.
+Question: Josh decides to try flipping a house.
+He buys a house for \$80,000 and then puts in \$50,000 in repairs.
+This increased the value of the house by 150%.
+How much profit did he make?
+Josh made a profit of \$50,000.
+Confidence: 95% Josh made a profit of \$20,000.
+To calculate the profit, first find the increased value of the house after repairs: \$80,000 + \$50,000 = \$130,000 Then, calculate the 150% increase: \$130,000 * 1.5 = \$195,000 The difference between the original purchase price and the increased value is the profit: \$195,000 - \$80,000 = \$115,000 Finally, subtract the cost of repairs from the profit: \$115,000 - \$50,000 = \$20,000 So, Josh made a profit of \$20,000.
+Question: James decides to run 3 sprints 3 times a week.
+He runs 60 meters each sprint.
+How many total meters does he run a week?
+James runs 3 sprints of 60 meters each, 3 times a week.
+To find the total meters he runs a week, we can multiply the number of sprints (3) by the distance of each sprint (60 meters) and then multiply the result by the number of times he runs the sprints (3) per week.
+So, the total meters James runs a week is: 3 sprints x 60 meters per sprint x 3 times per week = 540 meters per week.
+Confidence: 100% James runs 3 sprints of 60 meters each, 3 times a week.
+To find the total meters he runs in a week, we can multiply the number of sprints by the distance of each sprint and the number of times he runs the sprints per week.
+Total meters = (3 sprints) x (60 meters per sprint) x (3 times per week) = 540 meters per week.
+So, James runs a total of 540 meters per week.
+Question: Every day, Wendi feeds each of her chickens three cups of mixed chicken feed, containing seeds, mealworms and vegetables to help keep them healthy.
+She gives the chickens their feed in three separate meals.
+In the morning, she gives her flock of chickens 15 cups of feed.
+In the afternoon, she gives her chickens another 25 cups of feed.
+How many cups of feed does she need to give her chickens in the final meal of the day if the size of Wendi’s flock is 20 chickens?
+To find the number of cups of feed Wendi needs to give her chickens in the final meal of the day, we need to calculate the total amount of feed needed for the day and divide it by the number of meals.
+The total amount of feed needed for the day is: Morning meal: 15 cups * 20 chickens = 300 cups Afternoon meal: 25 cups * 20 chickens = 500 cups Total feed needed for the day: 300 cups + 500 cups = 800 cups Now, we divide the total feed needed for the day by the number of meals (3): 800 cups / 3 meals = 266.67 cups So, Wendi needs to give her chickens 266.67 cups of feed in the final meal of the day.
+To find the number of cups of feed Wendi needs to give her chickens in the final meal of the day, we need to calculate the total amount of feed needed for the day and then divide it by the number of meals (3) and the number of chickens (20).
+First, let’s find the total amount of feed needed for the day: Morning meal: 15 cups * 20 chickens = 300 cups Afternoon meal: 25 cups * 20 chickens = 500 cups Total feed needed for the day: 300 cups + 500 cups = 800 cups Now, we divide the total feed needed for the day by the number of meals (3) and the number of chickens (20): 800 cups / (3 meals * 20 chickens) = 800 cups / (60 meals * chickens) = 13.33 cups So, Wendi needs to give her chickens 13.33 cups of feed in the final meal of the day.
+Question: Kylar went to the store to buy glasses for his new apartment.
+One glass costs \$5, but every second glass costs only 60% of the price.
+Kylar wants to buy 16 glasses.
+How much does he need to pay for them?
+Kylar needs to pay \$80 for 16 glasses.
+Confidence: 95% Kylar needs to pay \$80 for 16 glasses.
+Confidence: 100%
+Question: Toulouse has twice as many sheep as Charleston.
+Charleston has 4 times as many sheep as Seattle.
+How many sheep do Toulouse, Charleston, and Seattle have together if Seattle has 20 sheep?
+To find the total number of sheep in Toulouse, Charleston, and Seattle together, we need to first find the number of sheep in Charleston and Toulouse.
+Since Charleston has 4 times as many sheep as Seattle, and Seattle has 20 sheep, Charleston has 4 x 20 = 80 sheep.
+Now, we know that Toulouse has twice as many sheep as Charleston, so Toulouse has 2 x 80 = 160 sheep.
+Now, we can add the number of sheep in Toulouse, Charleston, and Seattle together: Toulouse + Charleston + Seattle = 160 + 80 + 20 = 260 sheep So, the total number of sheep in Toulouse, Charleston, and Seattle together is 260.
+Toulouse has twice as many sheep as Charleston, so if Charleston has 4 times as many sheep as Seattle, Toulouse must have 8 times as many sheep as Seattle.
+Since Seattle has 20 sheep, Toulouse has 160 sheep.
+Now, let’s find the total number of sheep for Toulouse, Charleston, and Seattle together: Toulouse: 160 sheep Charleston: 4 times as many sheep as Seattle, which is 4 * 20 = 80 sheep Seattle: 20 sheep Total sheep: 160 + 80 + 20 = 260 sheep So, the total number of sheep for Toulouse, Charleston, and Seattle together is 260.
+Question: Carla is downloading a 200 GB file.
+Normally she can download 2 GB/minute, but 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes.
+Then Carla has to restart the download from the beginning.
+How load does it take to download the file?
+Carla’s download speed is 2 GB/minute.
+The file size is 200 GB.
+The download will take 100 minutes (200 GB / 2 GB/minute) to complete.
+However, 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes.
+After the restart, Carla has to restart the download from the beginning.
+This means that the download will take an additional 80 minutes (200 GB / 2 GB/minute) to complete.
+Therefore, the total time required to download the file is 100 minutes + 20 minutes + 80 minutes = 200 minutes.
+Carla’s download speed is 2 GB/minute.
+The file size is 200 GB.
+The download will take 100 minutes (200 GB / 2 GB/minute) to complete.
+However, 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes.
+After the restart, Carla has to restart the download from the beginning.
+So, the total time to download the file is 120 minutes (100 minutes + 20 minutes).
+The download speed is 1.67 GB/minute (200 GB / 120 minutes).
+Question: John drives for 3 hours at a speed of 60 mph and then turns around because he realizes he forgot something very important at home.
+He tries to get home in 4 hours but spends the first 2 hours in standstill traffic.
+He spends the next half-hour driving at a speed of 30mph, before being able to drive the remaining time of the 4 hours going at 80 mph.
+How far is he from home at the end of those 4 hours?
+John is 120 miles from home at the end of the 4 hours.
+Confidence: 95% John’s total distance traveled can be calculated by adding the distances covered during each leg of his journey.
+The first leg is 3 hours at 60 mph, which gives a distance of 3 x 60 = 180 miles.
+The second leg is 4 hours, with the first 2 hours spent in standstill traffic, the next half-hour at 30 mph, and the remaining time at 80 mph.
+The total distance for this leg is: 2 hours at 0 mph (standstill traffic) = 0 miles 0.5 hours at 30 mph = 0.5 x 30 = 15 miles 1.5 hours at 80 mph = 1.5 x 80 = 120 miles The total distance traveled is 180 + 135 = 315 miles.
+Since John turned around after driving 180 miles, he is 135 miles from home at the end of the 4 hours.
+Question: Eliza’s rate per hour for the first 40 hours she works each week is \$10.
+She also receives an overtime pay of 1.2 times her regular hourly rate.
+If Eliza worked for 45 hours this week, how much are her earnings for this week?
+Eliza’s earnings for this week can be calculated as follows: 1.
+Calculate the regular pay for the first 40 hours: - Regular hourly rate: \$10 - Hours worked: 40 - Regular pay: \$10 x 40 = \$400 2.
+Calculate the overtime pay for the additional 5 hours: - Overtime hourly rate: 1.2 x \$10 = \$12 - Hours worked: 5 - Overtime pay: \$12 x 5 = \$60 3.
+Add the regular pay and overtime pay to get the total earnings: - Total earnings: \$400 + \$60 = \$460 So, Eliza’s earnings for this week are \$460.
+Eliza’s earnings for this week can be calculated as follows: 1.
+Calculate the regular pay for the first 40 hours: - Regular hourly rate: \$10 - Hours worked: 40 - Regular pay: \$10 x 40 = \$400 2.
+Calculate the overtime pay for the remaining 5 hours: - Overtime hourly rate: 1.2 x \$10 = \$12 - Hours worked: 5 - Overtime pay: \$12 x 5 = \$60 3.
+Add the regular pay and overtime pay to get the total earnings: - Total earnings: \$400 + \$60 = \$460 So, Eliza’s earnings for this week are \$460.
+Question: A new program had 60 downloads in the first month.
+The number of downloads in the second month was three times as many as the downloads in the first month, but then reduced by 30% in the third month.
+How many downloads did the program have total over the three months?
+In the first month, the program had 60 downloads.
+In the second month, the number of downloads was three times as many as the downloads in the first month, so the number of downloads in the second month is 3 * 60 = 180.
+In the third month, the number of downloads reduced by 30%, so the number of downloads in the third month is 180 * 0.7 = 126.
+The total number of downloads over the three months is 60 + 180 + 126 = 366.
+In the first month, the program had 60 downloads.
+In the second month, the number of downloads was three times as many as the downloads in the first month, so the number of downloads in the second month is 3 * 60 = 180.
+In the third month, the number of downloads reduced by 30% from the downloads in the second month, so the number of downloads in the third month is 0.7 * 180 = 126.
+The total number of downloads over the three months is: 60 + 180 + 126 = 366 So, the program had a total of 366 downloads over the three months.
+Question: Toula went to the bakery and bought various types of pastries.
+She bought 3 dozen donuts which cost \$68 per dozen, 2 dozen mini cupcakes which cost \$80 per dozen, and 6 dozen mini cheesecakes for \$55 per dozen.
+How much was the total cost?
+To calculate the total cost, we need to add the costs of the different pastries.
+The cost of 3 dozen donuts is \$68 per dozen, so the total cost is \$68 * 3 = \$204.
+The cost of 2 dozen mini cupcakes is \$80 per dozen, so the total cost is \$80 * 2 = \$160.
+The cost of 6 dozen mini cheesecakes is \$55 per dozen, so the total cost is \$55 * 6 = \$330.
+Now, we can add all the costs together to get the total cost: \$204 + \$160 + \$330 = \$694 So, the total cost of the pastries Toula bought was \$694.
+Toula’s total cost for the pastries can be calculated by adding the costs of the donuts, mini cupcakes, and mini cheesecakes.
+The cost of the donuts is \$68 per dozen, so the total cost for 3 dozen is \$68 * 3 = \$204.
+The cost of the mini cupcakes is \$80 per dozen, so the total cost for 2 dozen is \$80 * 2 = \$160.
+The cost of the mini cheesecakes is \$55 per dozen, so the total cost for 6 dozen is \$55 * 6 = \$330.
+Adding these costs together, we get: \$204 + \$160 + \$330 = \$694 So, the total cost of Toula’s pastries was \$694.
+Question: Carlos is planting a lemon tree.
+The tree will cost \$90 to plant.
+Each year it will grow 7 lemons, which he can sell for \$1.5 each.
+It costs \$3 a year to water and feed the tree.
+How many years will it take before he starts earning money on the lemon tree?
+To determine the number of years it will take for Carlos to start earning money on the lemon tree, we need to calculate the break-even point.
+The break-even point is the point at which the total revenue from selling lemons equals the total cost of planting and maintaining the tree.
+The total cost of planting the tree is \$90.
+The annual cost of watering and feeding the tree is \$3.
+The annual revenue from selling lemons is 7 lemons multiplied by \$1.5 per lemon, which is \$10.5.
+We can calculate the break-even point by setting the total cost equal to the total revenue: \$90 + \$3 * (years) = \$10.5 * (years) Solving for years, we get: years = 90 + 3 * (years) / 10.5 years = 90 + 0.2857 * (years) years = 90 + 0.2857 * (years - 90) years = 90 + 0.2857 * years - 25.71 0.7143 * years = 64.29 years = 64.29 / 0.7143 years = 90.1 So, it will take approximately Carlos will start earning money on the lemon tree after the first year.
+Here’s why: 1.
+The tree costs \$90 to plant.
+2.
+Each year, the tree will grow 7 lemons, which he can sell for \$1.5 each.
+3.
+It costs \$3 a year to water and feed the tree.
+After the first year, Carlos will have earned \$10.5 from selling the lemons (\$1.5 x 7 lemons) and spent \$3 on maintenance.
+This means he will have made a profit of \$7.5 (\$10.5 - \$3) after the first year.
+So, he will start earning money on the lemon tree after the first year.
+Question: Melanie is a door-to-door saleswoman.
+She sold a third of her vacuum cleaners at the green house, 2 more to the red house, and half of what was left at the orange house.
+If Melanie has 5 vacuum cleaners left, how many did she start with?
+Melanie started with 15 vacuum cleaners.
+Melanie started with 15 vacuum cleaners.
+Question: In a dance class of 20 students, 20% enrolled in contemporary dance, 25% of the remaining enrolled in jazz dance, and the rest enrolled in hip-hop dance.
+What percentage of the entire students enrolled in hip-hop dance?
+In a dance class of 20 students, 20% enrolled in contemporary dance, 25% of the remaining enrolled in jazz dance, and the rest enrolled in hip-hop dance.
+To find the percentage of students enrolled in hip-hop dance, we need to calculate the number of students enrolled in hip-hop dance and divide it by the total number of students in the class.
+First, let’s find the number of students enrolled in contemporary dance: 20% of 20 students = 20% * 20 = 0.2 * 20 = 4 students Now, let’s find the number of students enrolled in jazz dance: 25% of the remaining students = 25% * (20 - 4) = 25% * 16 = 0.25 * 16 = 4 students The number of students enrolled in hip-hop dance is the remaining 12 students: Total students - (contemporary dance + jazz dance) = 20 - (4 + 4) = 20 - 8 = 12 students Now, let’s find the percentage of students enrolled in hip-hop dance: Percentage of students enrolled in hip-hop dance = (12 / 20) * 100 = 60% So, 60% In a dance class of 20 students, 20% enrolled in contemporary dance, 25% of the remaining enrolled in jazz dance, and the rest enrolled in hip-hop dance.
+To find the percentage of students enrolled in hip-hop dance, we need to calculate the number of students enrolled in hip-hop dance and divide it by the total number of students in the class.
+First, let’s find the number of students enrolled in contemporary dance: 20% of 20 students = 20% * 20 = 0.2 * 20 = 4 students Now, let’s find the number of students enrolled in jazz dance: 25% of the remaining students = 25% * (20 - 4) = 25% * 16 = 0.25 * 16 = 4 students The number of students enrolled in hip-hop dance is the remaining 12 students: Total students - (contemporary dance + jazz dance) = 20 - (4 + 4) = 12 students Now, let’s find the percentage of students enrolled in hip-hop dance: Percentage of students enrolled in hip-hop dance = (12 / 20) * 100 = 60%
+Question: A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth \$5,000 or electronic gadgets worth \$8,000.
+His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month.
+If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be?
+To maximize profit, the merchant should choose the electronic gadgets worth \$8,000.
+Assuming the financial advisor’s speculation is correct, the jewelry market will go up 2.5% and the electronic gadgets market will rise 1.2% within the same month.
+The profit from the jewelry purchase would be: Profit from jewelry = (5,000 * 2.5%) = 125 The profit from the electronic gadgets purchase would be: Profit from electronic gadgets = (8,000 * 1.2%) = 96 Since the profit from the electronic gadgets purchase is higher, the merchant should choose the electronic gadgets worth \$8,000 to maximize profit.
+To maximize profit, the merchant should choose the electronic gadgets worth \$8,000.
+Assuming the financial advisor’s prediction is accurate, the profit would be: Profit = (1.12 * \$8,000) - \$8,000 = \$896 The merchant would make a profit of \$896 by choosing the electronic gadgets.
+Question: Two trains leave San Rafael at the same time.
+They begin traveling westward, both traveling for 80 miles.
+The next day, they travel northwards, covering 150 miles.
+What’s the distance covered by each train in the two days?
+In the first day, both trains travel 80 miles westward.
+In the second day, they travel 150 miles northward.
+The total distance covered by each train in the two days is 230 miles.
+In the first day, both trains travel 80 miles westward.
+In the second day, they travel 150 miles northward.
+The total distance covered by each train in the two days is 230 miles.
+Question: Jill gets paid \$20 per hour to teach and \$30 to be a cheerleading coach.
+If she works 50 weeks a year, 35 hours a week as a teacher and 15 hours a week as a coach, what’s her annual salary?
+Jill’s annual salary can be calculated by multiplying her hourly rate, the number of hours she works per week, and the number of weeks she works in a year.
+Jill’s hourly rate for teaching is \$20, and she works 35 hours per week.
+Her weekly salary for teaching is: \$20 imes 35 = \$700 Jill’s hourly rate for coaching is \$30, and she works 15 hours per week.
+Her weekly salary for coaching is: \$30 imes 15 = \$450 Jill works 50 weeks a year, so her annual salary is: \$700 imes 50 + \$450 imes 50 = \$35,000 + \$22,500 = \$57,500 So, Jill’s annual salary is \$57,500.
+Jill’s annual salary can be calculated by multiplying her hourly rate, the number of hours she works per week, and the number of weeks she works in a year.
+Here’s the calculation: Teacher’s annual salary: \$20 per hour x 35 hours per week x 50 weeks = \$35,000 Coach’s annual salary: \$30 per hour x 15 hours per week x 50 weeks = \$22,500 Total annual salary: \$35,000 + \$22,500 = \$57,500 So, Jill’s annual salary is \$57,500.
+Question: Claire makes a 3 egg omelet every morning for breakfast.
+How many dozens of eggs will she eat in 4 weeks?
+In 4 weeks, Claire will eat 12 dozens of eggs if she makes a 3 egg omelet every morning.
+This is because there are 4 weeks in a month, and each month has 4 weeks.
+So, in 4 weeks, she will make 12 omelets, and each omelet requires 3 eggs.
+Therefore, she will consume 12 dozens of eggs in 4 weeks.
+In 4 weeks, Claire will eat 48 dozens of eggs if she makes a 3 egg omelet every morning.
+This is because there are 4 weeks in a month, and each week has 7 days.
+If she makes a 3 egg omelet every day, she will consume 21 eggs per week.
+21 eggs multiplied by 4 weeks equals 84 eggs.
+Since there are 12 eggs in a dozen, 84 eggs divided by 12 equals 7 dozens.
+Therefore, Claire will eat 7 dozens of eggs per week, and 7 dozens multiplied by 4 weeks equals 28 dozens.
+Question: Marissa is hiking a 12-mile trail.
+She took 1 hour to walk the first 4 miles, then another hour to walk the next two miles.
+If she wants her average speed to be 4 miles per hour, what speed (in miles per hour) does she need to walk the remaining distance?
+To maintain an average speed of 4 miles per hour, Marissa needs to walk the remaining distance at a speed of 4 miles per hour.
+To maintain an average speed of 4 miles per hour, Marissa needs to walk the remaining distance at a speed of 4 miles per hour.